### v1

# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

Abstract
Steering vision foundation models at inference time without retraining or access to large labeled datasets is a desirable yet challenging objective, particularly in dynamic or resource-constrained settings. In this paper, we introduce Visual Sparse Steering (VS2), a lightweight, test-time method that guides vision models using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 surpasses zero-shot CLIP by 4.12% on CIFAR-100, 1.08% on CUB-200, and 1.84% on Tiny-ImageNet. We further propose VS2++, a retrieval-augmented variant that selectively amplifies relevant sparse features using pseudo-labeled neighbors at inference time. With oracle positive/negative sets, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to better align the sparse features learned through the SAE reconstruction task with those relevant for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS). By incorporating a prototype-alignment loss during SAE training, using labels only during training while remaining fully test-time unsupervised, PASS consistently, though modestly, outperforms VS2, achieving a 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
1 Introduction
Foundation Models (FMs) have exhibited strong generalization capabilities across a wide range of domains, from classification tasks to open-ended generation  clip ; sam ; llama2 ; llava ; gpt4 . Despite their impressive performance, these models largely function as black boxes, restricting their controllability, and reliability in practice. The standard approach to edit a model’s behavior involves first collecting a labeled dataset resembling the target objective, followed by a supervised fine-tuning (SFT) procedure upon the holistic zhang2023instruction ; dong2023abilities or partial (e.g., PEFT lester2021power ; lora ; ia3 ) parameter space. Nevertheless, this procedure is cumbersome and can become prohibitively expensive given the model scale. Consequently, Steering Vector (SV) methods turner2023activation ; zou2023representation ; liu2024incontext , as an alternative, have gained popularity in the Large Language Model (LLM) domain as they can effectively modulate the model’s generation at inference-time in a training-free manner. While SVs have shown promise in LLMs, their functionality in the vision domain remains unclear.
In this work, we investigate the functionality of steering vectors in vision foundation models (VFMs). In particular, we use zero-shot classification tasks as a testbed to explore how SVs can be leveraged to upweight relevant visual features. Conventional SVs operate by first constructing a directional vector within the latent space to indicate a desired change of the functionality. It then linearly combines the directional vector with the original latent embeddings to craft output embeddings. In LLMs, SVs rely on two symbolic anchors that are inherently interpretable to determine the functional direction of the SV. For example, a happiness SV can be extracted using the latent embedding arithmetic corresponding to Happy - Sad. The formulation of SVs in LLM poses two critical challenges in vision domain. Given an image at inference-time: (1) Unlike symbolic language tokens, there exists tremendous redundancy in visual representations; consequently, it is unclear what features should be steered; (2) Without inherently interpretable anchors (e.g., happy, sad), towards which direction to steer is not obvious.
(a) VS2: Visual Sparse Steering w/o Contrastive Data
(b) VS2++: Selective Visual Sparse Steering
Figure 1: Overview of VS2 and VS2++. At inference, VS2 finds sparse latent concepts, and steers the original embedding towards the direction of amplifying those sparse features. When additional caching data is available, VS2 selectively enhances certain features while suppressing others in a contrastive manner.
To address the aforementioned challenges, we propose Visual Sparse Steering (VS2). Drawing inspiration from the interpretability provided by sparse auto-encoders (SAEs) hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , we hypothesize that the sparse features in a SAE’s latent space can capture key aspects of a visual embedding. VS2 upweights these features equally to construct a steering vector for classification. This equal upweighting of features is imprecise in that we expect different features to be more relevant than others for classification tasks. Despite this, we show that the simple equal upweighting of features still improves zero-shot classification.
To construct a more precise steering vector, we propose VS2++, which leverages additional unlabeled data to construct steering vectors for classification. For a given target image, VS2++ retrieves the Top-N most similar visual embeddings, and constructs positive and negative groups according to generated pseudo labels [^1] In practice, pseudo labels can be gathered using either unsupervised clustering methods or zero-shot classification results. The VS2++ SV is then constructed via subtracting the average SAE features of the negative group from the positive group. Intuitively, VS2++ upweights the SAE features most relevant for distinguishing the positive and negative groups. In summary, VS2 and VS2++ aim to enhance downstream classification performance while retaining the interpretability of SAEs. To further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS).
VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP  radford2021learning by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively, without any external data or supervision. When using external unlabeled image data, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Finally, by adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Our main contributions are:
  • We investigate the functionality of steering vectors in vision domain, and propose VS2 that leverages sparse feature learning to construct steering vectors for zero-shot classification.
  • We further propose VS2++ for scenarios where additional unlabeled data is available to selectively amplify important concepts. In this case, we use VS2 on pseudo-positive and -negative data to upweight relevant features for classification.
  • We propose Prototype-Aligned Sparse Steering (PASS), a lightweight method that introduces a prototype-alignment loss during SAE training using class labels only during training while remaining fully test-time unsupervised.
  • We empirically demonstrate that VS2, VS2++, and PASS can leverage SAEs to substantially improve zero-shot classification performance, outperforming baseline methods by a large margin under their corresponding regimes, while also retaining interpretability.
2 Related Work
Mechanistic Interpretability and Sparse Autoencoders.
Several traditional approaches exist for interpretability in vision models, including feature visualizationsimonyan2014visualising ; zeiler2014visualizing ; olah2017feature and network dissectionbau2017network ; oikarinen2022clip . Mechanistic interpretability seeks to systematically analyze and understand neural networks elhage2021mathematical ; olah2020zoom , but it faces challenges due to polysemantic neurons i.e. units that activate in response to multiple, seemingly unrelated inputs elhage2022toy . This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations elhage2022toy . Sparse Autoencoders (SAEs) have been shown to mitigate superposition by applying sparse dictionary learning to model internals sharkey2022taking ; bricken2023monosemanticity . Recent efforts have leveraged SAEs to uncover interpretable features within LLMs, revealing latent units tied to grammar rules, style patterns, and factual knowledge templeton2024scaling ; cunningham2023sparse ; gao2024scaling .  joshi2025identifiable refines this by training autoencoders on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision. While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
While some studies investigate SAEs for exploratory feature analysis or generative modeling in the visual domain hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , they primarily address concept discovery rather than performance enhancement. Other attempts in the vision domain try to create new SAE architectures for concept extraction  fel2025archetypal ; thasarathan2025universal . In contrast, our method leverages SAEs not only to interpret but also to actively steer vision models, yielding classification improvements alongside interpretable latent representations.
Similar to our line of work, joseph2025steering validate CLIP’s steerability and its efficacy in suppressing typographic attacks. Patch-SAElim2024sparse improves zero-shot classification by applying latent masking based on class-level activations. Specifically, it aggregates SAE activations from the training split to identify the top-$k$ most frequently activated latents per class, which are then used to mask active latents at test time. Our approach differs in a key aspect: we do not rely on activation aggregations from the same class in training but our method is instead fully test-time unsupervised. This design avoids reliance on class-conditioned statistics, making our approach more flexible and applicable to a wide range of tasks without requiring dataset-specific activation distributions. Specifically, our method (i) operates entirely in the vision latent space, (ii) requires no class labels, external vocabularies, or gradient updates, and (iii) improves zero-shot classification by amplifying sparse SAE features rather than replacing or suppressing activations.
Steering Vectors.
Steering vector (SV) methods turner2023activation ; park2023the ; hernandez2024linearity ; NIPS2013_9aa42b31 , also known as representation engineering zou2023representation , construct a directional vector (or task vector) and apply it in the latent space to change the target model’s behavior at inference-time. SVs have gained popularity in LLM/MLLM domain as it can improve generation security liu2024incontext , promoting truthfulness NEURIPS2023_81b83900 , mitigating hallucination vista , as well as improving inference efficiency liu2024incontext ; i2cl . Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts gavrikov2025can . In contrast, we differentiate from this line of work by focusing on Vision Foundation Models directly steering representations in the latent space without any external language signal.
Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions bayat2025steering ; makelov2024sparse . Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled concept pairs, making it more adaptable to general visual representations. Furthermore, its applicability to vision foundation models (VFM) remains unclear. The proposed visual sparse steering (VS2) method demonstrates that SV methods, combined with sparse autoencoders, can be utilized to enhance image features, improving zero-shot classification.
3 Method
3.1 Preliminaries
Sparse Autoencoder (SAE). A SAE consists of an encoder-decoder pair that maps an input vector ${\bm{x}}$ into a latent representation ${\bm{z}}$, then reconstructs ${\bm{x}}$ from ${\bm{z}}$. Formally, we define the encoder as:
${\bm{z}}=\mathrm{ReLU}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}\Bigr{)},$ (1)
where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input. The decoder then reconstructs ${\bm{x}}$ as:
$\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}},$ (2)
where ${\bm{W}}_{\mathrm{dec}}\in{\mathbb{R}}^{d\times n}$ is the decoder weight matrix. To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective is:
$\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ (3)
In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  gao2024scaling explicitly enforces the sparsity by using a top-$k$ selection mechanism:
${\bm{z}}=\mathrm{TopK}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})\Bigr{)},$ (4)
where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest. This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
3.2 Visual Sparse Steering
In this section, we introduce Visual Sparse Steering (VS2), a novel framework that leverages sparse features to steer neural representations toward meaningful semantic directions. Our approach targets two interconnected objectives, i.e., (1) explicitly identifying semantically meaningful concepts embedded within latent representations and (2) utilizing these interpretable concepts to define steering directions that enhance downstream task performance by emphasizing relevant semantic information and reducing spurious correlations. Unlike standard steering methods that rely on explicitly curated positive and negative examples, our approach utilizes Sparse Autoencoders (SAEs) to automatically discover and isolate sparse concepts that assist in defining a steering direction. By selectively amplifying these concepts within the latent space, VS2 generates conceptually grounded steering vectors that improve model interpretability, enhance classification accuracy, and robustly mitigate reliance on superficial or irrelevant cues.
3.2.1 VS2: Visual Sparse Steering
Learning latent concept space. Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding ${\bm{x}}$ into a sparse concept activation vector:
${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|}).$ (5)
By imposing sparsity constraints described in Sec. 3.1 on the concept activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful concepts. To map the concepts into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these concepts:
$\tilde{{\bm{x}}}=D_{c}({\bm{c}}).$ (6)
Since explicit supervision for concept identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful concepts.
Constructing steering vector. At inference stage, emerged sparse concepts are deemed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$. In implementation, we simply scale the original concept activations ${\bm{c}}$ by a factor of $\gamma$ to amplify their effects:
${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}.$ (7)
Reconstructing this modified concept vector ${\bm{c}}^{\prime}$ provides a conceptually steered embedding:
$\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime}).$ (8)
Since reconstruction steps inherently introduce approximation errors, we define the steering vector ${\bm{v}}$ as:
${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}},$ (9)
which explicitly represents the semantic shift induced by upweighting and mitigates the effects of the reconstruction errors.
Steer the target emebedding. Obtaining the steering vector ${\bm{v}}$ that aims to amplify the latent concepts embedded in ${\bm{x}}$, we shift the original embedding by adding the steering vector back as:
$\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}.$ (10)
where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale:
$\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ (11)
Intuitively, this procedure precisely moves the embedding ${\bm{x}}$ along directions defined by semantically relevant, interpretable concepts, thus improving interpretability and enhancing robustness against spurious correlations. We refer Fig. 1 (upper) for a schematic overview of VS2.
3.2.2 VS2++: Visual Sparse Steering with Cached Data
In VS2, we leverage the learned SAEs to implicitly decide the critical concepts for a given latent embedding, and treat all identified concepts equally important. One natural speculation is that there are certain concepts that are unique and more important to the target underlying distribution, while others are more general even spuriously correlated. To this end, we propose VS2++ that allows the access of cached training embedding at test-time to enable a more nuanced control over the concept steering.
Constructing Contrastive Groups. Given a query embedding ${\bm{x}}_{q}$ and a cached dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
$\mathcal{N}_{N}({\bm{x}}_{q})\;=\;\underset{\begin{subarray}{c}\{{\bm{x}}_{i}\}\subseteq\{{\bm{x}}_{i}\}_{i=1}^{M}\\ |\mathcal{N}_{N}({\bm{x}})|=N\end{subarray}}{\mathrm{arg\,max}}\;S({\bm{x}}_{q},{\bm{x}}_{i}),$ (12)
where $S(,)$ denotes the measurement of cosine similarity.
After retrieving $\mathcal{N}_{N}({\bm{x}}_{q})$, we obtain the pseudo-label for the query embedding $\hat{y}_{q}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{q})$, where $P(y\,\mid\,{\bm{x}})$ is the classifier’s predicted probability for label $y$. Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$.
We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding:
${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ (13)
Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors:
${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$ (14)
Selective Steering. For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first applies VS2 (Eq. 7,8,9) to acquire a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying concepts within the given embedding. We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$. We then average these per-embedding steering vectors within each group, yielding the positive anchor
$\bar{{\bm{v}}}^{p}\;=\;\frac{1}{\bigl{|}{\mathbb{S}}^{+}({\bm{x}}_{q})\bigr{|}}\,\sum_{{\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})}{\bm{v}}_{i}^{p},$ (15)
and the negative anchor
$\bar{{\bm{v}}}^{n}\;=\;\frac{1}{\bigl{|}{\mathbb{S}}^{-}({\bm{x}}_{q})\bigr{|}}\,\sum_{{\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})}{\bm{v}}_{j}^{n}.$ (16)
Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor:
${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ (17)
and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
$\hat{{\bm{x}}}_{q}\;=\;{\bm{x}}_{q}\;+\;\lambda\,{\bm{v}},\quad\hat{{\bm{x}}}_{q}\;=\;\frac{\hat{{\bm{x}}}_{q}\,\cdot\,\|{\bm{x}}_{q}\|_{2}}{\|\hat{{\bm{x}}}_{q}\|_{2}}.$ (18)
This selective scheme thus anchors the steering directions in concepts consistently activated within the positive group, while suppressing undesired features highlighted by the negative group. Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases.
4 Experiments
Table 1: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and Without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Zero-shot (no retrieval) $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 ) RAG-enhanced (oracle cache) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle cache) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
Experimental Setup.
We test VS2 and VS2++ on CIFAR-100 krizhevsky2009learning , Tiny-ImageNet le2015tiny , and CUB-200 wah2011caltech three datasets, covering standard, complex, and fine-grained classification tasks. These datasets vary in scale, diversity, and class distribution, allowing us to assess VS2 and VS2++ in both general and imbalanced classification scenarios. For vision foundation model, we use CLIP radford2021learning with ViT-B/32, and ViT-B/16 backbones. We train top-$k$ Sparse Autoencoders (SAEs) on the CLS token embeddings extracted from every layer of CLIP’s vision transformer, using the training data of each dataset to learn sparse latent representations. More details regarding the training of the top-$k$ Sparse Autoencoders can be found in the Appendix C.
4.1 Visual Sparse Steering methods
4.1.1 VS2: Visual Sparse Steering without Contrastive Data
Baselines.
We evaluate the effectiveness of VS2 in the absence of external contrastive data. As baselines, we report: (1) the zero-shot performance of the original CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{CLIP}_{\text{REC}}^{\text{F}}$, which uses the SAE-reconstructed CLS token from the final layer without any steering; and (3) $\text{CLIP}_{\text{REC}}^{\text{A}}$, which uses the reconstructed CLS tokens from all layers. These baselines isolate the effect of reconstruction alone, without applying any modification or steering vector. All top-$k$ SAEs are trained on CLS token representations across all transformer layers. In addition to these baselines, we report $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ that applies steering by upweighting the top-$k$ sparse features of the reconstructed CLS token of the final layer with a scaling factor of $\gamma=1.5$. This variant isolates the impact of steering the SAE’s latent activations by amplifying their top-$k$ before the CLS embedding is reconstructed, letting us quantify the benefit attributable to this targeted up-weighting alone. In Appendix G, we provide the pseudocode for main differences between the proposed baselines and VS2. In Appendix H, we also provide comparative results of our method compared to other baselines such as Splice (bhalla2024interpreting, ). Note that Splice uses an external vocabulary to construct latent features; in contrast, we use unsupervised SAEs to learn latent features for VS2 and VS2++.
Results.
From Table 1, we demonstrate that using only SAE reconstructions generally reduces zero-shot performance in all datasets; we hypothesize this is due to reconstruction errors in SAEs (engels2025decomposing, ). However, when SAE features are amplified in the reconstruction, performance can improve, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ for CIFAR-100. Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% using the ViT-B/32 and ViT-B/16 backbones respectively. However, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ is worse than the CLIP baseline on CUB-200 and Tiny-ImageNet. In contrast, our method VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. In summary, we demonstrate that our proposed steering vectors using SAEs can improve downstream classification performance on vision tasks without relying on any external data or supervision. In Appendix A, we also provide evidence that the sparse features are interpretable; they can capture subtle class-specific attributes and concepts such as birds with gray upperpart, white underpart, white-colored throat and black eyes. Finally, in Appendix B, we provide a sensitivity analysis of VS2 hyperparameters $\gamma$ and $\lambda$ showing the robustness of our proposed methods.
4.1.2 VS2++: Visual Sparse Steering with Selectively Amplifying Features
In VS2, we leverage learned SAEs to identify critical concepts from a latent embedding, treating all activated concepts as equally important. However, some concepts may be more discriminative or unique to the target class, while others may be generic or even spuriously correlated with a class label. To enable finer control over concept importance, we introduce VS2++, which accesses cached training data at test time to guide steering more selectively. As a proxy for identifying salient features, we adopt a Retrieval-Augmented Generation (RAG) approach: we retrieve the top-$50$ most similar training images to the test input using DINOv2 oquab2023dinov2 and average their sparse activations. To construct the positive and negative sets as shown in equations 13, 14, we assign pseudolabels using a CLIP classifier; images with the most frequent pseudolabel form the positive set, while the remaining images serve as negatives. This setup corresponds to the Non-oracle cache in Table 1. For comparison, we also report an unrealistic scenario that serves as an upper-bound i.e. Oracle cache setting that uses ground-truth labels to define positives and negatives.
Baselines.
As a baseline in the Retrieval-Augmented Generation (RAG) pipeline lewis2020retrieval , we combine the test query embedding with a weighted aggregation of retrieved document embeddings. Additional details on the weighting strategy are provided in Appendix D. We also include a non-SAE variant that constructs a steering vector by averaging the CLIP embeddings of the positive and negative sets and taking their difference, yielding a purely contrastive direction without relying on Sparse Autoencoders.
Results.
In Table 1, we present the results for VS2 ++ when external data is available under both RAG-enhanced settings where we have oracle and non-oracle cache data. When we have oracle positive and negative sets, we observe that VS2++ surpasses all other methods under almost every setting. Weighted RAG performs worse than the steering vector approaches (both CLIP and VS2++), suggesting that directionality is more informative than proximity. In the more realistic scenario where we do not have oracle positive and negative sets, we observe drops in accuracies. We hypothesize that this degradation is due to inaccuracies in pseudolabeling, resulting in the amplification of spurious directions in the SAE latent space. However, across the six configurations, VS2++ outperforms the CLIP Steering Vector in four cases and falls behind only in the ViT-B/32 runs for CIFAR-100 and Tiny-ImageNet (with CIFAR-100 essentially on par). These results suggest that there remains significant room for improvement in actively selecting which sparse features to steer, especially under noisy or weakly supervised retrieval scenarios. In Appendix I, we conduct ablation studies on the top-$N$ retrieved images that indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
4.1.3 Fine-Grained Per-Class Accuracy Analysis
To understand where Visual Sparse Steering helps most, we compute per-class top-1 accuracy on CIFAR-100 with a ViT-B/32 backbone and report the ten largest improvements in Table 2. The results reveal a long-tailed pattern: VS2 lifts accuracy by up to 25 % in individual classes (e.g. lawn-mower $\rightarrow$ tractor, pine-tree $\rightarrow$ forest), while VS2++ reaches gains of 38 % when high-quality neighbours are available (e.g. bee $\rightarrow$ spider, lion $\rightarrow$ tiger, whale $\rightarrow$ flatfish). In all cases, the corrected predictions involve visually or taxonomically proximate categories, indicating that sparse steering primarily disambiguates fine-grained confusions rather than providing a uniform boost. This behaviour aligns with our qualitative findings in the Appendix A.2, where the top-$k$ SAE features capture subtle class-specific attributes (birds with gray upperpart, white underpart, white-colored throat and black eyes etc.) that are otherwise entangled in the original CLIP embedding. A more detailed list of per-class performance accuracies of VS2 and VS2++ is provided in Table 9 in the Appendix.
Table 2: Top-10 class gains on CIFAR-100. Numbers are top-1 accuracies; green = absolute gain over zero-shot (ZS).
Class ZS VS2 (+$\Delta$) Mislabel Why visually confusing tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower Both are wheeled agricultural machines; similar viewed side-on. forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree Dense tree canopies often dominated by tall conifers. man $53.0$ 74.0 $\uparrow$(+21.0) boy Male human silhouette; age difference subtle at CIFAR resolution. bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck Rectangular vehicle profile with windows and wheels. snake $61.0$ 76.0 $\uparrow$(+15.0) worm Long, legless bodies with smooth textures. woman $57.0$ 71.0 $\uparrow$(+14.0) girl Female human figures; height/size cues are small. trout $18.0$ 32.0 $\uparrow$(+14.0) aquarium_fish Similar fish shape and reflective scales in water scenes. cockroach $49.0$ 62.0 $\uparrow$(+13.0) beetle Small, dark exoskeleton insects with segmented bodies. bridge $72.0$ 83.0 $\uparrow$(+11.0) sea Large spans photographed above wide blue water areas. rose $58.0$ 68.0 $\uparrow$(+10.0) poppy Red, multi-petal flowers with central dark core.
(a) VS2
Class ZS VS2++ (+$\Delta$) Mislabel Why visually confusing spider $53.0$ 91.0 $\uparrow$(+38.0) bee Small multi-legged insects; dark radial silhouettes. tiger $48.0$ 84.0 $\uparrow$(+36.0) lion Large orange/brown big cats; stripes vs. mane often unclear at low resolution. flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale Flattened marine shapes against blue water backgrounds. possum $30.0$ 65.0 $\uparrow$(+35.0) hamster Small brown furry mammals with rounded bodies. lizard $39.0$ 74.0 $\uparrow$(+35.0) snake Slender reptiles; legs on lizard hard to spot in $32\times 32$ images. sweet_pepper $58.0$ 92.0 $\uparrow$(+34.0) orange Round orange-coloured produce with glossy skins. raccoon $49.0$ 83.0 $\uparrow$(+34.0) skunk Mid-sized nocturnal mammals with dark fur and contrasting markings. shrew $46.0$ 79.0 $\uparrow$(+33.0) skunk Small ground mammals; dark elongated silhouettes blur distinguishing cues. snail $53.0$ 86.0 $\uparrow$(+33.0) mushroom Brown dome (shell vs. cap) resting on the forest floor. beetle $49.0$ 82.0 $\uparrow$(+33.0) cockroach Dark, shiny exoskeletal insects with similar body shapes.
(b) VS2++
4.2 Prototype-Aware Sparse Steering Vectors
Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE). These directions correspond to latent concepts that, ideally, align with class-discriminative features. Steering in these directions upweights what the model has learned to be important during reconstruction. This raises a central hypothesis: the reconstruction task itself is to some extent sufficient to uncover features that are also relevant for classification. In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task. In this section, drawing inspiration from prototype theory rosch1973natural , we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
The limited improvements observed on fine-grained datasets like CUB-200 and Tiny-ImageNet suggest that the challenge lies not just in identifying sparse features, but in uncovering the correct ones. This shifts the central question from “What are the most important sparse features to select?” to a deeper inquiry: “Can the reconstruction objective alone reliably capture features that are most useful for classification and if not, how can task-relevant information be effectively incorporated?”
Oracle steering with known prototypes.
Following prototype theory, we collect for every class the ten images that ViT-B/32 CLIP classifies with the highest confidence; these serve as oracle prototypes. Given the true class label $y$, we build a prototype steering vector by averaging their latent sparse features. VS2 using oracle prototypes can lift CLIP to 97.5% on CIFAR-100, 91.04% on CUB-200, and 90.1% on Tiny-ImageNet. This confirms the existence of discriminative sparse directions. In Appendix E we further examine the discriminative ability (measured by orthogonality) of these steering vectors. Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
Prototype-aligned SAE (PASS).
Above, we constructed steering vectors using the pretrained SAE features of oracle prototypes. Now, we consider whether prototypes can be used to learn new, more informative SAE features. We assume during SAE training that we have access to class labels. Then, to the SAE loss we add a regularization term which encourages SAE features to be close to their class mean. That is, for a training sample $i$ with latent sparse feature  $\mathbf{z}_{i}$ and class mean $\bar{\mathbf{z}}_{\text{class}(i)}$ we minimize
$\mathcal{L}=\mathcal{L}_{\text{recon}}+w_{\text{aux}}\;\bigl{\lVert}\mathbf{z}_{i}-\bar{\mathbf{z}}_{\text{class}(i)}\bigr{\rVert}_{2}^{2},$ (19)
where $w_{\text{aux}}$ controls the strength of the prototype-alignment term relative to the reconstruction loss and is set to 0.8. We refer to the resulting steering method as PASS (Prototype-Aligned Sparse Steering). Although PASS uses class labels during SAE training, it remains fully test-time unsupervised. Empirically, in Table 1, we observe that PASS outperforms VS2 across all datasets, with particularly substantial gains on CIFAR-100. However, this improvement comes at the cost of requiring labels for each training sample during SAE training. Gains are modest on CUB-200 and Tiny-ImageNet, and we thus hypothesize that classes which share many features require richer or multi-prototype guidance which is an intriguing avenue for future work.
5 Conclusion
We introduce Visual Sparse Steering VS2 and its retrieval-augmented variant VS2++, two test-time methods that steer vision foundation models using sparse features obtained from top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively. To control concept importance more precisely, we propose VS2++. With oracle positive/negative sets, VS2++ surpasses every baseline in nearly every configuration, achieving absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. When those sets are noisy, its accuracy drops but still surpasses the standard Steering Vector in most cases, highlighting the need for better feature selection under weakly supervised retrieval. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we proposed Prototype-Aligned Sparse Steering (PASS). By adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Selecting appropriate prototypes or exploring multi-prototype guidance is an intriguing avenue for future work.
Appendix
Appendix A Decoding the Sparse Latent Space: Insights from SAEs
Our proposed methods, VS2 and VS2++, achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs). We hypothesize that SAEs identify meaningful sparse concepts, which in turn guide the steering mechanisms. To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned concepts.
A.1 Quantitative Evaluation of Concept Significance.
To further evaluate the role of interpretable sparse features, we manipulate the top-$k$ most active latent dimensions extracted via Sparse Autoencoders. This experiment examines whether these features are critical for classification and how model predictions change under different modifications. We explore the following manipulation settings:
  1. 1.
    Zeroing-Out ($\gamma=0$): We set the top-$k$ most active features to zero before applying the steering vector. This removes their influence while preserving the remaining latent structure.
  2. 2.
    Negation ($\gamma=-1$): We invert the sign of the top-$k$ features before applying the steering vector, effectively pushing the representation in the opposite direction. This tests whether these dimensions encode class-discriminative information.
Table 3 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention. We observe that negating or zeroing the top-$k$ most important concepts consistently degrades performance across all datasets. This result confirms that the concepts learned by the Sparse Autoencoders are essential for classification.
Table 3: Effect of manipulating top-$k$ sparse codes. Zero-shot accuracy (%) drops sharply when dominant sparse features are zeroed ($\gamma=0$) or negated ($\gamma=-1$), confirming their importance.
Modification CIFAR-100 CUB-200 Tiny-IN
$\text{CLIP}_{\text{ZS}}$ $61.07$ $38.68$ $56.64$
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $37.30$ $39.49$
+ Zero-out ($\gamma=0$) $1.71$ $1.14$ $16.11$
+ Negate ($\gamma=-1$) $0.06$ $0.00$ $0.82$
A.2 Qualitative Evaluation of Concept Significance.
We qualitatively investigate the concepts learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE). Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature. Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes. For the qualitative evaluation of concept significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations. Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images. Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension. The intuition is that if a specific concept frequently emerges among the top-activating images for a particular dimension, that dimension is strongly associated with that concept.
For feature 511 as shown in Figure 2 (left), the activated images exhibit consistent semantic characteristics, including a gray upper part and a white underpart. Notably, this feature predominantly activates for images from similar but different classes, specifically different types of Gulls, such as Western, California, Herring, and Slaty-Backed Gulls. We observe that the top-activating images for these classes share all concepts in common which is not the case with dimension 3067. For feature 3067, as shown in Figure 2 (right) we observe that the top-activating images share common visual attributes, such as a white-colored throat and black eyes. Through human qualitative evaluation, we find that features in the sparse latent space capture meaningful visual concepts, grouping semantically similar images together, either from the same class or across different classes, as long as they share underlying conceptual similarities.
Figure 2: Concept coverage analysis of learned sparse latent features in the Sparse Autoencoder (SAE). Each subfigure illustrates the top-activating images for two different SAE dimensions, highlighting the consistency of shared visual concepts among the highest-activated examples. The analysis demonstrates how sparse features capture meaningful semantic attributes, grouping semantically similar images either within or across classes. Left: Top-activating images for feature 511. The images predominantly belong to different classes of Gulls (e.g., Western, California, Herring, and Slaty-Backed Gulls), yet they share consistent visual characteristics such as a gray upper part and a white underpart. This suggests that feature 511 captures a semantically meaningful concept spanning multiple related categories. Right: Top-activating images for feature 3067. The images share distinct visual attributes, including a white-colored throat and black eyes. However, unlike feature 511, these images belong to more diverse categories, indicating that this latent dimension captures a broader concept that generalizes across different classes.
Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
We analyze how the two key hyperparameters in VS2 (i) the sparse-feature amplification $\gamma$, and (ii) the steering vector scale $\lambda$ affect downstream accuracy. In the absence of contrastive supervision, these parameters govern how strongly we amplify sparse activations and how far the embedding is shifted in feature space. We sweep both values across a grid on three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet using ViT-B/32 backbone.
Figure 3 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees. Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy. For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$. Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse codes and embedding distortion. The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
(a) CIFAR-100
(b) CUB-200
(c) Tiny-ImageNet
Figure 3: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$. All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$. Accuracy degrades if either parameter becomes too large.
Appendix C Training top-$k$ Sparse Autoencoders
We follow CLIP [18] with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token. Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer. We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned. We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units. Training largely follows [31] and uses  [53], with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
We monitor reconstruction quality with the fraction of variance unexplained (FVU; lower is better), defined as
$\text{FVU}\;=\;\frac{\lVert\mathbf{X}-\hat{\mathbf{X}}\rVert_{F}^{2}}{\lVert\mathbf{X}-\bar{\mathbf{X}}\rVert_{F}^{2}},$ (20)
where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance). FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 00 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean. In Table 4, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
Table 4: Fraction of variance unexplained (FVU; lower is better). Each SAE has 4× expansion; sparsity $k=64$ for ViT-B/32 and $k=256$ for ViT-B/16.
Dataset ViT-B/32 (k=64) ViT-B/16 (k=256)
CIFAR-100 $0.2812$ $0.1166$
CUB-200 $0.2487$ $0.1653$
Tiny-ImageNet $0.5060$ $0.3018$
Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline. We use DINOv2 [50] to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
$\mathbf{E}=\alpha\mathbf{q}+(1-\alpha)\sum_{j=1}^{k}w_{j}\mathbf{r}_{j},$
where $\mathbf{q}$ is the query embedding, $\mathbf{r}_{j}$ the $j$-th retrieved embedding, and $w_{j}$ the normalized similarity weight:
$w_{j}=\frac{s_{j}}{\sum_{i=1}^{k}s_{i}},\quad\text{where }s_{j}=\text{sim}(\mathbf{q},\mathbf{r}_{j}).$
The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set.
We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 4 shows results for CIFAR-100, and Tiny-ImageNet.
Findings.
CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance. On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result. This suggests that the retrieval set carries strong distributional signal in cluttered datasets. Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context.
The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines. For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ work best; for noisy domains, more reliance on the retrieval pool is beneficial.
(a) CIFAR-100
(b) Tiny-ImageNet
Figure 4: RAG sensitivity to $\alpha$ and top-$k$. Accuracy varies with the weight $\alpha$ on the original query and the number $k$ of retrieved images. Larger $k$ often introduces noise; smaller $\alpha$ performs better on cluttered datasets.
Appendix E Are exemplar-derived directions really distinct?
A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B. Using the oracle prototypes as described in Section 4.2, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder. Most pairs have low similarity ($\!\text{mean}\!=\!0.23$), yet a non-negligible tail reveals strongly overlapping directions. Table 5 lists the ten highest-overlap pairs:
Table 5: Top-10 most overlapping prototype steering directions on CIFAR-100. High cosine similarity indicates that the two classes share visual attributes that the SAE encodes along nearly the same latent axis.
Rank Class 1 Class 2 Cosine $\uparrow$
1 beetle cockroach $0.91$
2 mouse shrew $0.89$
3 dolphin shark $0.84$
4 otter seal $0.84$
5 dolphin whale $0.84$
6 possum raccoon $0.84$
7 snake worm $0.83$
8 oak tree willow tree $0.83$
9 ray shark $0.81$
10 bowl cup $0.80$
These high-overlap pairs are semantically plausible confusions (e.g. beetle vs. cockroach or dolphin vs. whale), confirming that exemplar steering directions tend to align for visually or taxonomically proximate classes. In downstream applications, a simple orthogonalization step may help reduce feature overlap between sparse directions. Investigating principled ways to encourage orthogonality during SAE training is a promising direction for future work.
Appendix F Ablation Study of Expansion factor and top-$k$
In Table 6, we present the downstream task accuracy of CLIP ViT-B/32 using various values of expansion factor and $k$. Across a 4 $\times$ range in width and an 8 $\times$ range in sparsity, top-1 accuracy fluctuates by less than one percentage point evidence that VS2’s performance is largely insensitive to the precise SAE capacity–sparsity trade-off. Additionally, in Table 7, we present the average cosine similarities of different steering vectors coming from various configurations of SAEs in terms of expansion factor and top-$k$.
Table 6: VS2 accuracy as a function of SAE width (expansion factor) and sparsity (top-$k$). Numbers are top-1 / top-5 (%). The best result is boldfaced; every other configuration is within of the optimum, highlighting the method’s robustness to architectural choices.
SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
4 $\times$, $k$=128 64.61 87.95
8 $\times$, $k$=128 64.56 87.76
4 $\times$, $k$=64 64.54 87.79
16 $\times$, $k$=64 64.54 87.78
8 $\times$, $k$=64 64.52 87.96
10 $\times$, $k$=128 64.43 87.81
16 $\times$, $k$=512 64.42 87.71
8 $\times$, $k$=512 64.40 87.91
4 $\times$, $k$=256 64.34 87.62
8 $\times$, $k$=256 64.29 87.80
16 $\times$, $k$=256 64.28 87.75
4 $\times$, $k$=512 64.12 87.87
16 $\times$, $k$=128 64.10 87.79
Table 7: Cosine similarity ($\uparrow$) between steering vectors learned by different SAE capacities. Rows/columns are ordered by expansion factor and sparsity (e × expansion, $k$ active).
SAE cfg.
16×/64
16×/512
16×/128
16×/256
8×/128
8×/512
8×/256
4×/256
4×/512
4×/128
10×/128
4×/64
8×/64
16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716
16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282
16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678
16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744
8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829
8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506
8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624
4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600
4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549
4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655
10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678
4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814
8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000
Appendix G Visual Sparse Steering Pseudocode
For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.1.
Algorithm 1 SAE_Steering – SAE-based CLS token modification
1:function SAE_Steering($\mathbf{h},\mathrm{SAE},k,\gamma,\lambda_{\Delta z},\text{mode}$) $\triangleright$ $\mathbf{h}$: CLS token; SAE: sparse autoencoder; $k$: sparsity level
2:     $\mathbf{a},\mathbf{idx}\leftarrow\mathrm{SAE.select\_topk}(\mathrm{SAE.pre\_acts}(\mathbf{h}),k)$
3:     $\mathbf{z}_{\text{base}}\leftarrow\mathrm{SAE.decode}(\mathbf{a},\mathbf{idx})$
4:     $\mathbf{z}_{\text{boost}}\leftarrow\mathrm{SAE.decode}(\gamma\cdot\mathbf{a},\mathbf{idx})$
5:     if $\text{mode}=\texttt{"reconstruction"}$" then
6:         return $\mathbf{z}_{\text{base}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
7:     else if $\text{mode}=\texttt{"amplified"}$" then
8:         return $\mathbf{z}_{\text{boost}}$triangleright Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
9:     else if $\text{mode}=\texttt{"steering"}$" then
10:         return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$triangleright VS2
11:     end if
12:end function
Appendix H Comparison with other Baseline Methods
We compare VS2 against SpLiCE [16], using the official implementation provided by the authors. For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration. Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
Table 8: Zero-shot top-1 accuracy (%) on CIFAR-100.
Method ViT-B/32 ViT-B/16 $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) SpLiCE [16] 55.57 ( –5.50 ) 58.29 ( –5.67 ) VS2 (ours) 64.52 ( +3.45 ) 68.08* ( +4.12 )
Appendix I Effect of Top-$N$ Retrieved Neighbors
To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation. We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet. As shown in Figure 5, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$. In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples. These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
Figure 5: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors $N$, using ViT-B/16 (Patch-16) across datasets. Larger $N$ improves general classification but may degrade performance in fine-grained settings.
Appendix J Per-class Performance Analysis on CIFAR-100
In Table 9, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++. We highlight both positive and negative shifts in accuracy to identify the most impacted categories. Unlike the average performance reported in the main experiments, this analysis shows that the top-10 class-level gains are generally larger in magnitude than the corresponding losses. Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible. As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
Table 9: Top-10 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15) woman $0.57$ 0.71 $\uparrow$(+0.14) trout $0.18$ 0.32 $\uparrow$(+0.14) cockroach $0.49$ 0.62 $\uparrow$(+0.13) bridge $0.72$ 0.83 $\uparrow$(+0.11) rose $0.58$ 0.68 $\uparrow$(+0.10)
(a) VS2 – Top-10 Gains
Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07) flatfish $0.55$ 0.49 (-0.06) willow_tree $0.46$ 0.41 (-0.05) lamp $0.75$ 0.70 (-0.05) lawn_mower $1.00$ 0.96 (-0.04) kangaroo $0.72$ 0.68 (-0.04)
(b) VS2 – Top-10 Losses
Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39) raccoon $0.44$ 0.83 $\uparrow$(+0.39) lizard $0.36$ 0.74 $\uparrow$(+0.38) wolf $0.54$ 0.90 $\uparrow$(+0.36) shrew $0.44$ 0.79 $\uparrow$(+0.35) bear $0.51$ 0.85 $\uparrow$(+0.34)
(c) VS2++ – Top-10 Gains
Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03) lawn_mower $1$ 0.98 (-0.02)
(d) VS2++ – Top-10 Losses