### v2

# Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors

Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors
Gerasimos Chatzoudis    Zhuowei Li    Gemma E. Moran    Hao Wang    Dimitris N. Metaxas
Machine Learning, ICML
Steering vision foundation models at test time, without updating foundation-model weights or using labeled target data, is a desirable yet challenging goal. We present Visual Sparse Steering (VS2), a lightweight, label-free adaptation method that constructs a steering vector from sparse features extracted by a Sparse Autoencoder (SAE) trained on unlabeled in-domain training-split activations of the vision encoder. VS2 offers three key advantages over existing test-time adaptation methods: (1) a feature-level intervention space in sparse SAE representations; (2) efficiency, requiring only a forward pass with no test-time optimization or backpropagation; and (3) a reliability diagnostic based on SAE reconstruction loss that can skip steering when reconstruction is poor, enabling safe fallback to the baseline, a capability not standard in conventional steering vectors and test-time adaptation methods. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves zero-shot top-1 accuracy by 3.45–4.12%, 0.93–1.08%, and 1.50–1.84%, respectively, while remaining forward-only and adding minimal compute overhead. A retrieval-based upper-bound analysis suggests substantial headroom if task-relevant sparse features can be selected reliably, motivating future work on selective feature amplification for interpretable, efficient test-time steering.
1 Introduction
Adapting vision foundation models at test time, without labeled data or weight updates, is desirable but challenging. Many test-time adaptation (TTA) methods rely on backpropagation through large encoders, incurring substantial compute overhead  (Shu et al., 2022; Feng et al., 2023; Yoon et al., 2024). Steering vectors offer a lightweight alternative but typically require carefully curated contrastive examples, limiting practicality in vision  (Turner et al., 2023; Zou et al., 2023; Liu et al., 2024). This motivates two challenges: (i) Can we construct effective steering directions without contrastive supervision or test-time optimization? and (ii) Can we find representations that mitigate the redundancy and entanglement inherent in visual features?
Drawing inspiration from mechanistic interpretability work on sparse autoencoders (SAEs) (Elhage et al., 2022; Bricken et al., 2023), we hypothesize that sparse SAE latents capture salient aspects of the input and can provide a less redundant basis for intervention in visual representations. SAEs have been used primarily as an interpretability tool to extract human-interpretable features in trained foundation models. However, their potential as a control mechanism remains underexplored: it is unclear whether sparse features can be manipulated in a principled, label-free manner to improve downstream performance. This gap motivates our approach: rather than using SAE features for post-hoc analysis, we leverage them as an actionable intervention space for efficient, label-free test-time steering.
To this end, we propose Visual Sparse Steering (VS2): we train a top-$k$ SAE on unlabeled in-domain activations and, at test time, construct a steering vector by amplifying the image’s active sparse features and projecting the induced change back to embedding space. VS2 is forward-only, requiring a single SAE forward pass with no test-time optimization, backpropagation, or contrastive anchor examples. Across CIFAR-100, CUB-200, and Tiny-ImageNet with two CLIP backbones (ViT-B/32, ViT-B/16)  (Radford et al., 2021b), VS2 improves zero-shot top-1 accuracy by 3.45%-4.12%, 0.93%-1.08%, and 1.50%-1.84%, respectively, while adding $0.1\%$ compute overhead.
A practical challenge is distribution shift: if a test input lies outside the SAE’s training distribution, reconstruction quality can degrade, and the resulting sparse features may be unreliable, causing steering to underperform. To mitigate this, VS2 includes a built-in reliability diagnostic based on SAE reconstruction error. When reconstruction error is high, we skip steering and fall back to the original zero-shot CLIP prediction, avoiding harmful adaptation from misaligned sparse features. This provides a simple, model-internal mechanism for deciding when to apply steering, which is not typically available in standard steering-vector or test-time adaptation methods.
A natural question is whether all active sparse features should be amplified uniformly, or whether selectively amplifying task-relevant features could yield larger gains. To quantify this headroom, we introduce VS2++ as a retrieval-based upper-bound analysis: given an external unlabeled corpus, it retrieves nearest neighbors, forms positive and negative groups using CLIP pseudo-labels, and constructs a contrastive steering direction in SAE feature space. With oracle positives/negatives, VS2++ yields gains of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate and motivating future work on robust selection under noisy pseudo-labels.
In summary, our contributions are:
  • Unsupervised sparse steering for vision. We introduce VS2, a lightweight, label-free method that trains a top-$k$ SAE on unlabeled in-domain activations and performs forward-only test-time steering without updating CLIP weights or using contrastive anchor examples.
  • Consistent gains with minimal overhead. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 improves zero-shot top-1 accuracy while adding $0.1\%$s compute overhead and requiring no test-time optimization.
  • Reliability-aware fallback. We propose a reconstruction-error-based diagnostic that can conservatively disable steering and fall back to the baseline when the SAE representation is unreliable, reducing the risk of harmful interventions under distribution shift, at the cost of occasionally forgoing potential gains under a strict reconstruction-error threshold.
  • Headroom via selective amplification. We present VS2++ as a retrieval-based upper-bound analysis showing substantial potential gains when task-relevant sparse features can be selected accurately, motivating future work on robust feature selection under noisy pseudo-labels.
(a) VS2: Visual Sparse Steering w/o Contrastive Data
(b) VS2++: Selective Visual Sparse Steering
Figure 1: Overview of VS2 and VS2++: at inference time, VS2 constructs a steering vector by equally amplifying sparse features, while VS2++ selectively enhances them when an external visual corpus is available.
2 Related Work
Mechanistic Interpretability and Sparse Autoencoders.
Several traditional approaches exist for interpretability in vision models, including feature visualization (Simonyan et al., 2014; Zeiler and Fergus, 2014; Olah et al., 2017) and network dissection (Bau et al., 2017; Oikarinen and Weng, 2022). Mechanistic interpretability seeks to systematically analyze and understand neural networks (Elhage et al., 2021; Olah et al., 2020), but it faces challenges due to polysemantic neurons i.e. units that activate in response to multiple, seemingly unrelated inputs (Elhage et al., 2022). This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations (Elhage et al., 2022). Sparse Autoencoders (SAEs) have been explored to mitigate superposition by applying sparse dictionary learning to model internals (Bricken et al., 2023). Recent efforts have leveraged SAEs to uncover interpretable features within LLMs  (Templeton, 2024; Cunningham et al., 2023; Gao et al., 2024a).  Joshi et al. (2025) train SAEs on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision. While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
While SAEs have been explored for feature analysis, generative modeling, and concept disentanglement in the visual domain (Bhalla et al., 2024; Stevens et al., 2025; Surkov et al., 2024; Fel et al., 2025; Thasarathan et al., 2025), these works focus primarily on interpretability rather than downstream performance. In contrast, our method leverages SAEs to actively steer vision models in a label-free, test-time setting to improve classification. Related efforts like (Joseph et al., 2025) study CLIP’s steerability on typographic attacks, while Patch-SAE (Lim et al., 2024) improves classification accuracy via class-conditioned latent masking. Our approach, on the other hand, requires no class labels and avoids class-based activation aggregation during training, enabling broader applicability without reliance on external supervision or gradient updates.
Steering vectors.
Steering Vector (SV) methods (Turner et al., 2023; Park et al., 2023; Hernandez et al., 2024; Mikolov et al., 2013), also known as representation engineering (Zou et al., 2023), construct a directional task vector and apply it in the latent space to change the target model’s behavior at inference time. In LLMs/MLLMs, SVs are used to enhance security (Liu et al., 2024), truthfulness (Li et al., 2023), reduce hallucinations (Li et al., 2025a), and improve efficiency (Li et al., 2025b). Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts (Gavrikov et al., 2025). In contrast, we focus on steering latent representations in vision models without any language input. Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions (Bayat et al., 2025; Makelov, 2024). Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled positive/negative concept pairs, making it more adaptable to general visual representations.
Test-Time Adaptation.
Recent advances in improving the generalization of vision-language models have introduced a variety of prompt-tuning approaches, such as CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), and MaPLe (Khattak et al., 2023), along with adapter-based techniques like Tip-Adapter (Zhang et al., 2021) and CLIP-Adapter (Gao et al., 2024b). Despite their effectiveness, these methods typically assume access to a small number of labeled target-domain examples (often in a few-shot setting). In contrast, Test-Time Adaptation methods aim to improve robustness under distribution shift by adapting a pre-trained model using only unlabeled test samples at inference time (Liu et al., 2021; Sun et al., 2020; Wang et al., 2020; Gao et al., 2022). Specifically, Test-Time Prompt Tuning (TPT) (Shu et al., 2022) introduced entropy-minimizing prompt optimization over augmented views, inspiring subsequent methods such as DiffTPT (Feng et al., 2023), which leverages diffusion-based augmentations, and C-TPT (Yoon et al., 2024), which incorporates calibration-aware objectives. Although effective, these strategies typically require optimization loops, multiple augmentations, and backpropagation through large encoders, resulting in substantial computational and memory overhead at inference. In contrast, VS2 performs a single forward pass with sparse SAE-guided steering, avoiding test-time optimization while still providing measurable gains.
3 Method
Preliminaries.
A Sparse Autoencoder (SAE) consists of an encoder-decoder pair that maps an input vector $\bm{x}$} into a latent representation $\bm{z}$}, then reconstructs $\bm{x}$} from $\bm{z}$}. Formally, we define the encoder as: ${\bm{z}}=\mathrm{ReLU}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}),$ where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input. The decoder then reconstructs $\bm{x}$} using a decoder weight matrix , as $\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}}$ . To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective becomes $\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  (Gao et al., 2024a) explicitly enforces the sparsity by using a top-$k$ selection mechanism: ${\bm{z}}=\mathrm{TopK}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})),$ where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest. This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
3.1 Visual Sparse Steering
Visual Sparse Steering (VS2) is a lightweight, label-free, test-time method that guides vision models using steering vectors derived from sparse features learned by top-k Sparse Autoencoders without requiring contrastive data. We hypothesize that the sparse features in an SAE’s latent space can capture salient aspects of the input, thereby reducing redundancy. To this end, VS2 constructs a steering vector by amplifying the SAE’s salient features, reconstructing the corresponding amplified representation, and subtracting the original reconstruction (without amplification) to minimize the inherent SAE’s reconstruction errors. The resulting sparse steering vector points in the direction of the salient features and is linearly combined with the original embedding to steer the model’s behavior. Unlike common steering vectors, VS2 does not require positive and negative anchor examples; instead, steering occurs in the direction of the most salient features identified by the SAE.
3.1.1 VS2: Visual Sparse Steering
Learning latent concept space. Given an input image $X\in\mathcal{X}$ and its corresponding embedding ${\bm{x}}\in{\mathbb{R}}^{d}$ extracted from layer $i$ of a Vision Foundation Model (VFM), we define a Concept Encoder $E_{c}\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{|{\mathbb{C}}|}$ that maps the embedding $\bm{x}$} into a sparse feature activation vector: ${\bm{c}}=E_{c}({\bm{x}})=(c_{1},\dots,c_{|{\mathbb{C}}|})$. By imposing sparsity constraints on the feature activations, the encoder $E_{c}$ is encouraged to identify disentangled, semantically meaningful features. To map the features into the original representation space, we define a Concept Decoder $D_{c}\colon{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{d}$ that reconstructs the original embedding from these features: $\tilde{{\bm{x}}}=D_{c}({\bm{c}})$. Since explicit supervision for feature identification is typically unavailable, we train the encoder-decoder pair jointly on training data $\mathcal{X}$, enforcing both sparsity and accurate embedding reconstruction to implicitly uncover meaningful features.
Constructing steering vector. At inference time, emerged sparse features are assumed semantically significant and critical for downstream tasks, and we enhance them via a Concept Upweighting Function $U:{\mathbb{R}}^{|{\mathbb{C}}|}\to{\mathbb{R}}^{|{\mathbb{C}}|}$. In implementation, we simply scale the original feature activations $\bm{c}$} by a factor of $\gamma$ to amplify their effects: ${\bm{c}}^{\prime}=U({\bm{c}})=\gamma\times{\bm{c}}$. Reconstructing this modified feature vector $\bm{c}$$^{\prime}$ provides a conceptually steered embedding: $\tilde{{\bm{x}}}^{\prime}=D_{c}({\bm{c}}^{\prime})$. Since reconstruction steps inherently introduce approximation errors, we define the steering vector $\bm{v}$} as: ${\bm{v}}=\tilde{{\bm{x}}}^{\prime}-\tilde{{\bm{x}}}$. which explicitly represents the semantic shift induced by upweighting and aims to mitigate the effects of the reconstruction errors.
Steer the target embedding. Obtaining the steering vector $\bm{v}$} that aims to amplify the latent features embedded in $\bm{x}$}, we shift the original embedding by adding the steering vector back as: $\hat{{\bm{x}}}={\bm{x}}+\lambda\,{\bm{v}}$. where $\lambda$ is a hyperparameter that controls the magnitude of steering. To stabilize the representation, we further rescale the $\ell_{2}$ norm of the steered embedding to its original scale: $\hat{{\bm{x}}}=\frac{\hat{{\bm{x}}}\cdot\|{\bm{x}}\|_{2}}{\|\hat{{\bm{x}}}\|_{2}}.$ Intuitively, this procedure precisely moves the embedding $\bm{x}$} along directions defined by semantically relevant, salient features, thus enhancing robustness against spurious correlations. We refer Fig. 1 (left) for a schematic overview of VS2.
3.1.2 VS2++: Visual Sparse Steering with Unlabeled Data
In VS2, we leverage the learned SAEs to identify the most salient features for a given latent representation, but treat all identified features equally important by equally amplifying them. In vision models, there are potentially many redundant features, e.g., when predicting labels, we want to ignore background features such as the sky. To selectively amplify only important features more robustly and non-uniformly, we relax the problem, allowing the existence of an unlabeled external vision corpus, and propose VS2++ that leverages this additional unlabeled data to construct steering vectors. For a given target image, VS2++ retrieves the Top-N most similar visual embeddings and constructs positive and negative groups according to generated pseudo labels. The VS2++ Steering Vector is then constructed via subtracting the average SAE features of the negative group from the positive group. Intuitively, VS2++ amplifies the SAE features most relevant for distinguishing the positive and negative groups.
Constructing Contrastive Groups. Given a query embedding ${\bm{x}}_{q}$ and an unlabeled dataset of embeddings $\{{\bm{x}}_{i}\}_{i=1}^{M}$, VS2++ first retrieve the set of $N$ nearest neighbors $\mathcal{N}_{N}({\bm{x}}_{q})$ for a query embedding ${\bm{x}}_{q}$ is defined as:
$\mathcal{N}_{N}({\bm{x}}_{q})\;=\;\underset{\begin{subarray}{c}\{{\bm{x}}_{i}\}\subseteq\{{\bm{x}}_{i}\}_{i=1}^{M}\\ |\mathcal{N}_{N}({\bm{x}})|=N\end{subarray}}{\mathrm{arg\,max}}\;S({\bm{x}}_{q},{\bm{x}}_{i}),$ (1)
where $S(,)$ denotes the measurement of cosine similarity. After retrieving $\mathcal{N}_{N}({\bm{x}}_{q})$, we obtain the pseudo-label for the query embedding $\hat{y}_{q}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{q})$, where $P(y\,\mid\,{\bm{x}})$ is the classifier’s predicted probability for label $y$. Similarly, for each neighbor ${\bm{x}}_{i}\in\mathcal{N}_{N}({\bm{x}}_{q})$, we compute its pseudo-label: $\hat{y}_{i}\;=\;\arg\max_{y}\,P(y\,\mid\,{\bm{x}}_{i})$. We then construct the positive group, ${\mathbb{S}}^{+}({\bm{x}}_{q})$, by selecting neighbors sharing the same pseudo-label as the query embedding: ${\mathbb{S}}^{+}({\bm{x}}_{q})\;=\;\{{\bm{x}}_{i}\,\in\,\mathcal{N}_{N}({\bm{x}}_{q})\;\mid\;\hat{y}_{i}=\hat{y}_{q}\}.$ Likewise, the negative group, ${\mathbb{S}}^{-}({\bm{x}}_{q})$, is formed by the remaining neighbors: ${\mathbb{S}}^{-}({\bm{x}}_{q})\;=\;\mathcal{N}_{N}({\bm{x}}_{q})\;\setminus\;{\mathbb{S}}^{+}({\bm{x}}_{q}).$
Selective Steering. For each embedding ${\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})$, VS2++ first obtains the steering vectors as in VS2 acquiring a directional vector ${\bm{v}}_{i}^{p}$ that underscores the underlying features within the given embedding. We repeat the same procedure for each ${\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})$ to obtain the negative steering vectors ${\bm{v}}_{j}^{n}$. We then average these per-embedding steering vectors within each group, yielding the positive anchor
$\bar{{\bm{v}}}^{p}\;=\;\frac{1}{\bigl|{\mathbb{S}}^{+}({\bm{x}}_{q})\bigr|}\,\sum_{{\bm{x}}_{i}\in{\mathbb{S}}^{+}({\bm{x}}_{q})}{\bm{v}}_{i}^{p},$ (2)
and the negative anchor
$\bar{{\bm{v}}}^{n}\;=\;\frac{1}{\bigl|{\mathbb{S}}^{-}({\bm{x}}_{q})\bigr|}\,\sum_{{\bm{x}}_{j}\in{\mathbb{S}}^{-}({\bm{x}}_{q})}{\bm{v}}_{j}^{n}.$ (3)
Finally, we obtain the contrastive steering vector by subtracting the negative anchor from the positive anchor: ${\bm{v}}\;=\;\bar{{\bm{v}}}^{p}\;-\;\bar{{\bm{v}}}^{n},$ and apply it to the query embedding ${\bm{x}}_{q}$ in the same manner as VS2:
$\hat{{\bm{x}}}_{q}\;=\;{\bm{x}}_{q}\;+\;\lambda\,{\bm{v}},\quad\hat{{\bm{x}}}_{q}\;=\;\frac{\hat{{\bm{x}}}_{q}\,\cdot\,\|{\bm{x}}_{q}\|_{2}}{\|\hat{{\bm{x}}}_{q}\|_{2}}.$ (4)
This selective scheme thus anchors the steering directions in features consistently activated within the positive group, while suppressing undesired features highlighted by the negative group. Since VS2++ retrieves N-nearest embeddings of a query to construct both positive and negative groups, negative group contains non-trivial hard cases. We refer Fig. 1 (right) for a schematic overview of VS2++.
4 Experiments
Table 1: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
CIFAR-100 CUB-200 Tiny-IN Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 Label-free, unsupervised steering $\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 ) $\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( -3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( -26.08 ) 52.96 ( -8.12 ) $\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( -2.85 ) 63.42 ( -0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( -20.31 ) 54.84 ( -6.24 ) $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( -17.15 ) 58.81 ( -2.27 ) VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 ) RAG-enhanced (oracle unlabeled external data) Weighted RAG 76.43 ( +15.36 ) 69.78 ( +5.82 ) 58.32 ( +6.56 ) 60.65 ( +5.59 ) 72.53 ( +15.89 ) 67.84 ( +6.75 ) CLIP Steering Vector 81.85 ( +20.78 ) 84.12 ( +20.16 ) 56.42 ( +4.66 ) 61.51 ( +6.45 ) 80.38 ( +23.74 ) 84.07 ( +22.99 ) VS2++ 81.95 ( +20.88 ) 85.40 ( +21.44 ) 58.84 ( +7.08 ) 61.91 ( +6.85 ) 73.27 ( +16.63 ) 81.55 ( +20.47 ) RAG-enhanced (non-oracle unlabeled external data) CLIP Steering Vector 77.22 ( +16.15 ) 78.97 ( +15.01 ) 47.89 ( -3.87 ) 53.95 ( -1.11 ) 74.06 ( +17.42 ) 76.79 ( +15.71 ) VS2++ 77.11 ( +16.04 ) 79.14 ( +15.18 ) 52.81 ( +1.05 ) 57.02 ( +1.96 ) 72.12 ( +15.48 ) 76.92 ( +15.84 )
Table 2: TTA comparisons.
Method CIFAR CUB Tiny
Zero-shot 61.07 51.76 56.64
Ensemble 63.66 51.54 61.39
TPT 64.09 51.83 62.77
C-TPT 64.86 52.54 63.20
Diff-TPT 63.04 52.80 61.60
VS2 64.52 52.69 58.14
VS2 + Ens 65.48 52.47 62.79
VS2 + Ens + Aug 65.80 51.54 62.59
Table 3: Inference-time compute.
Method GFLOPs GMACs Params (M)
Plain CLIP (vision) 8.7295 4.3623 87.456
LoRA (r=16r{=}16) 8.7885 4.3918 88.046
VS2 (SAE steering) 8.7342 4.3647 92.178
We evaluate VS2 on three datasets: CIFAR-100 (Krizhevsky et al., 2009), Tiny-ImageNet (Le and Yang, 2015), and CUB-200 (Wah et al., 2011), covering standard, complex, and fine-grained classification tasks. For the vision foundation model, we use CLIP (Radford et al., 2021b) with ViT-B/32, and ViT-B/16 backbones. Our goal is to investigate the effectiveness of Sparse Autoencoders (SAEs) for constructing steering vectors. In Sections 4.1 and 4.2, we focus on the case where the SAE can faithfully reconstruct the test inputs. To this end, we train top-$k$ SAEs on CLS token embeddings extracted from each layer of CLIP’s vision transformer, using the training split of each dataset to learn sparse latent representations (unsupervised in-domain data). Additional training details are provided in Appendix C. In Section 4.4, we relax this assumption by training a single, more general SAE across all train datasets, including both general and fine-grained tasks, to better approximate a real-world deployment scenario and evaluate how VS2 can avoid harmful steering when test samples cannot be faithfully reconstructed. This capability is novel compared to common steering vector and test-time adaptation techniques.
(a) CIFAR-100
(b) CUB-200
(c) Tiny-ImageNet
Figure 2: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$. All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$. Accuracy degrades if either parameter becomes too large.
4.1 VS2: Visual Sparse Steering
Baselines.
We evaluate VS2 on downstream image classification and compare against baselines that isolate reconstruction and latent amplification effects. We report: (1) the zero-shot CLIP model ($\text{CLIP}_{\text{ZS}}$); (2) $\text{SAE}_{\text{REC}}^{\text{F}}$, which replaces the final-layer [CLS] token with its SAE reconstruction (no steering); and (3) $\text{SAE}_{\text{REC}}^{\text{A}}$, which reconstructs [CLS] tokens from all layers (no steering). These baselines measure the impact of reconstruction alone. We additionally include $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$, which scales the top-$k$ latent activations by a fixed factor ($\gamma=1.5$) before reconstructing the final-layer [CLS] token, serving as an ablation for latent amplification without our additive steering direction. Appendix G provides pseudocode for these variants. We also report comparisons to Splice (Bhalla et al., 2024) in Appendix H, which defines latents using an external vocabulary rather than learning an unsupervised SAE dictionary. Unless stated otherwise, VS2 applies steering to the final-layer [CLS] token; Appendix E analyzes steering across layers.
Results.
From Table 1, we show that using only SAE reconstructions typically reduces zero-shot performance; we attribute this degradation to reconstruction error in SAEs (Engels et al., 2025). However, when SAE features are amplified during reconstruction, performance can improve in some cases, as demonstrated by $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ on CIFAR-100. Specifically, $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ outperforms the zero-shot baseline on CIFAR-100 by 1.62% and 2.85% with ViT-B/32 and ViT-B/16, respectively, but underperforms CLIP on CUB-200 and Tiny-ImageNet. In contrast, VS2 consistently improves over zero-shot CLIP on all three datasets and both backbones, yielding gains of 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet for ViT-B/32 and ViT-B/16, respectively. These results demonstrate that sparsity-guided steering outperforms both reconstruction-only and amplification-only baselines. Although our focus is on how to steer ViTs using SAE features rather than interpretability, SAEs are often studied in mechanistic interpretability to identify human-interpretable features. Accordingly, Appendix A provides qualitative examples to assess whether the learned sparse features are consistent with subtle class-specific visual attributes (e.g., gray upperparts, white underparts, white throat, and black eyes in bird images). Finally, Figure 2 and Appendix B report a sensitivity analysis over the steering hyperparameters $\gamma$ and $\lambda$, showing a broad region of near-optimal performance and supporting the robustness of VS2 to reasonable hyperparameter choices.
Comparison with Test-Time Adaptation Baselines.
Table 2 compares VS2 to test-time prompt tuning methods TPT (Shu et al., 2022), C-TPT (Yoon et al., 2024), and Diff-TPT (Feng et al., 2023) using ViT-B/32. These baselines adapt by optimizing text prompts at test time, whereas VS2 performs no test-time optimization and does not update prompts or model weights. The Ensemble setting uses 80 fixed ImageNet-style templates following CLIP’s standard zero-shot protocol (Radford et al., 2021a). Aug denotes test-time image augmentations. The TTA baselines follow the standard 63-view protocol, while our VS2+Ens+Aug ablation uses 10 augmented views per image. Since the compared methods use different adaptation mechanisms and test-time compute, we treat this as an accuracy reference under common TTA evaluation protocols rather than a matched-compute comparison. Under this evaluation, VS2 matches or exceeds prompt-tuning baselines on CIFAR-100 and remains competitive on CUB-200; moreover, combining VS2 with prompt ensembling closes much of the gap on Tiny-ImageNet (VS2+Ens: 62.79% vs. TPT: 62.77%).
Computational Overhead.
VS2 is designed to be a lightweight alternative to parameter- or prompt-based test-time adaptation. Table 3 reports inference-time cost for the CLIP ViT-B/32 vision encoder under different adaptation strategies. Adding VS2 increases vision-encoder compute by only 0.0047 GFLOPs ($0.1\%$), since the SAE operates only on the [CLS] token at a single layer. In contrast, a LoRA adapter with rank 16 is applied inside every transformer block, leading to higher FLOPs despite fewer trainable parameters. At training time, learning the SAE is also inexpensive. With cached activations, SAE training requires only 0.014 GFLOPs per sample, compared to roughly 26 GFLOPs per sample for LoRA or full fine-tuning of the vision encoder; full training-time FLOPs and parameter counts are reported in Appendix M.
Finally, prompt-optimization-based TTA can be substantially more expensive at inference because it evaluates many augmented views and performs gradient-based updates. A single CLIP forward pass on a $224\times 224$ image with 100 text prompts costs approximately 0.59 TFLOPs (vision+text; Appendix M). Under the standard TPT protocol (63 augmented views and 4 optimization steps with a backward pass through the text encoder), this increases to roughly 9.8 TFLOPs per test image, i.e., $\approx 16.6\times$ the cost of plain CLIP (Appendix M). In contrast, VS2 adds only one SAE forward pass and keeps the per-image vision-encoder overhead below $0.1\%$ of CLIP.
4.2 VS2++: Upper-Bound Analysis via Selective Amplification
In VS2, we uniformly amplify the active sparse features of the test image. A natural extension is selective amplification: upweight features that are predictive for the target class while suppressing nuisance factors. To quantify the headroom available from accurate feature selection, we introduce VS2++ as a retrieval-based upper-bound analysis that uses neighbors to induce weak supervision for selecting which sparse features to steer.
Concretely, for each test image we retrieve the top-$50$ nearest neighbors from the dataset’s unlabeled training split using DINOv2 (Oquab et al., 2023). We then form pseudo-positive and pseudo-negative groups using CLIP predictions: neighbors with the most frequent pseudo-label constitute positives, and the remaining neighbors are treated as negatives. We construct a contrastive steering direction in SAE feature space by subtracting the average negative sparse activations from the average positive sparse activations. This corresponds to the non-oracle setting in Table 1. For comparison, we also report an oracle upper bound where positives and negatives are defined using the ground-truth labels of the retrieved neighbors.
We emphasize that VS2++ is used primarily to study the gains possible from reliable feature selection, motivating future work on robust selection under noisy pseudo-labels. As baselines, we include a weighted Retrieval Augmented Generation (RAG) pipeline (Lewis et al., 2020), which combines the query embedding with a weighted average of retrieved CLIP embeddings (Appendix D), as well as a non-SAE contrastive steering baseline that computes the difference between the mean CLIP embeddings of the positive and negative groups.
Results.
Table 1 reports VS2++ under oracle (upper bound) and non-oracle (CLIP pseudo-labeled) positive/negative sets. In the oracle setting, VS2++ achieves gains of up to 21.44% (CIFAR-100), 7.08% (CUB-200), and 20.47% (Tiny-ImageNet), demonstrating substantial headroom when feature selection is accurate. Weighted RAG is generally weaker than contrastive steering on CIFAR-100 and Tiny-ImageNet while remaining competitive on CUB-200, indicating that directional offsets are often more informative than proximity alone; VS2++ is competitive with CLIP-space steering but not uniformly best (e.g., Tiny-ImageNet with ViT-B/32). In the non-oracle setting, accuracy drops due to pseudo-label noise, yet VS2++ outperforms the CLIP steering vector in four of six configurations and does not exhibit negative transfer on CUB-200. Overall, these results motivate robust feature selection under weak supervision; Appendix I ablates the number of retrieved neighbors.
Table 4: Top-5 class gains on CIFAR-100. Top-1 accuracy; green = gain over CLIP zero-shot.
Class ZS VS2 / VS2++ Mislabel Why visually confusing VS2 tractor $55.0$ 80.0 $\uparrow$(+25.0) lawn_mower wheeled farm machines; side view forest $35.0$ 58.0 $\uparrow$(+23.0) pine_tree dense conifer canopy man $53.0$ 74.0 $\uparrow$(+21.0) boy age cue weak at $32{\times}32$ bus $51.0$ 68.0 $\uparrow$(+17.0) pickup_truck boxy vehicle silhouette snake $61.0$ 76.0 $\uparrow$(+15.0) worm elongated legless body VS2++ spider $53.0$ 91.0 $\uparrow$(+38.0) bee small dark radial silhouette tiger $48.0$ 84.0 $\uparrow$(+36.0) lion stripes vs. mane unclear flatfish $49.0$ 85.0 $\uparrow$(+36.0) whale flat marine shape on blue possum $30.0$ 65.0 $\uparrow$(+35.0) hamster small brown furry mammal lizard $39.0$ 74.0 $\uparrow$(+35.0) snake legs often invisible
4.3 Fine-Grained Per-Class Accuracy Analysis
[…]
4.4 SAEs as a reliability diagnostic for safe visual sparse steering
Our main experiments assume access to a sparse autoencoder whose reconstructions of the vision encoder activations are sufficiently faithful on the target distribution. We approximate this by training top-$k$ SAEs on unlabeled in-domain training-split activations for each dataset, which allows us to isolate and investigate the effectiveness of sparsity-guided steering vectors in modulating the classification behavior of CLIP’s vision transformer.
Relaxation: a generalized SAE.
In practice, a single high-fidelity SAE that remains faithful across diverse domains is difficult to obtain. To approximate this setting, we train a generalized SAE on the union of the unlabeled training splits of CIFAR-100, CUB-200, and Tiny-ImageNet and then apply this same SAE across datasets. Table 5 reports top-1 accuracy together with Fraction of Variance Unexplained (FVU) for generalized steering. Larger FVU values indicate poorer reconstruction, and FVU >1>1 implies worse reconstruction than a mean predictor. We observe that generalized steering improves over the zero-shot baseline on CIFAR-100 and Tiny-ImageNet, where reconstruction error is low (FVU $\approx 0.21$ on CIFAR-100 and $\approx 0.44$ on Tiny-ImageNet), but degrades on CUB-200, where reconstruction error is substantially higher (FVU $\approx 1.93$). This pattern suggests that steering effectiveness is sensitive to SAE fidelity, motivating the use of reconstruction error as a reliability signal.
Table 5: Reconstruction and classification with a generalized SAE. We report top-1 accuracy (%) with corresponding reconstruction error FVU in parentheses.
CIFAR-100 CUB-200 Tiny-IN
Method B/32 B/16 B/32 B/16 B/32 B/16
Baseline 61.07 (–) 63.96 (–) 51.76 (–) 55.06 (–) 56.64 (–) 61.08 (–)
VS2 64.63 (0.216) 68.22 (0.208) 48.81 (1.931) 49.22 (1.930) 59.73 (0.448) 63.83 (0.437)
Table 6: FVU-calibrated fallback for VS2 on in-distribution and out-of-distribution (OOD) datasets using a generalized SAE. Each cell reports top-1 accuracy with coverage (fraction steered) in parentheses.
In Distribution Out of Distribution
Method CIFAR-100 CUB-200 Tiny-IN Aircraft Food101 Flower102 Caltech101 SUN EuroSAT
[…]
VS2 (In Domain SAE, ungated) 64.52 52.69 58.14 27.69 81.20 64.81 80.95 54.80 34.65
VS2 (Generalized SAE, ungated) 64.63 48.81 59.73 18.57 81.06 64.11 79.72 45.78 30.76
VS2 + FVU gate (q=0.90q{=}0.90) 64.27 (99.93%) 51.76 (0.0%) 59.29 (85.0%) 24.87 (0.00%) 80.72 (0.00%) 63.70 (0.11%) 78.44 (2.30%) 51.80 (0.11%) 30.39 (54.20%)
VS2 + FVU gate (q=0.95q{=}0.95) 64.62 (99.97%) 51.76 (0.04%) 58.44 (92.16%) 24.87 (0.00%) 80.71 (0.02%) 63.72 (1.32%) 78.52 (5.51%) 51.80 (0.49%) 31.39 (63.15%)
VS2 + FVU gate (q=0.99q{=}0.99) 64.63 (100%) 51.79 (0.62%) 58.65 (98.63%) 24.87 (0.09%) 80.70 (0.89%) 64.34 (22.72%) 78.76 (23.82%) 51.92 (6.97%) 31.57 (85.52%)
VS2 + FVU gate (q=0.995q{=}0.995) 64.63 (100%) 51.83 (1.54%) 58.64 (99.34%) 24.87 (0.21%) 80.69 (3.10%) 64.21 (40.62%) 78.78 (35.26%) 52.11 (15.28%) 31.65 (92.09%)
Out-of-distribution steering with FVU-gated fallback.
We next evaluate robustness under distribution shift by applying the generalized SAE to six unseen datasets: Aircraft (Maji et al., 2013), Food101 (Bossard et al., 2014), Flower102 (Nilsback and Zisserman, 2008), Caltech101 (Fei-Fei, 2004), SUN (Xiao et al., 2010), and EuroSAT (Helber et al., 2019). As a reference point, Table 6 shows that when an in-domain SAE is used, VS2 consistently improves over zero-shot CLIP across all six OOD datasets. However, ungated generalized steering can substantially degrade performance on some OOD datasets (e.g., Aircraft and SUN), consistent with unreliable sparse features when reconstruction quality is poor. We therefore use per-sample FVU as a reliability diagnostic: if a test sample’s FVU exceeds a threshold $\tau$, we skip steering and fall back to the baseline CLIP embedding; otherwise, we apply VS2.
To set $\tau$ without labels, we calibrate a single threshold on the unlabeled CIFAR-100+Tiny-ImageNet training union by taking the $q$-quantile of per-sample FVU scores, and then apply the same $\tau$ across all datasets. For FVU-gated rows in Table 6, we report top-1 accuracy together with coverage, the fraction of samples for which steering is applied. Overall, FVU gating acts as a conservative safety mechanism: on datasets where generalized steering is harmful, coverage drops and performance moves toward the baseline, while on in-distribution datasets with reliable reconstruction, coverage remains high and the gains of generalized steering are largely retained.
5 Discussion and Future Work
We introduced VS2, a lightweight and fully label-free adaptation method that steers a frozen vision encoder using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders (SAEs). Unlike conventional steering vectors that rely on curated contrastive anchors, VS2 constructs a per-instance direction directly from the SAE’s sparse support: it amplifies the active latents, projects the induced change back to embedding space, and applies the resulting offset at inference time. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 consistently improves zero-shot accuracy while remaining forward-only and adding negligible compute overhead. More broadly, our results support the view that SAE latents can serve as an actionable intervention space for improving downstream behavior, beyond post-hoc interpretability.
A key practical constraint is SAE fidelity. When reconstruction is poor, the resulting sparse features can be unreliable and steering may become harmful. A central advantage of VS2 is that reconstruction error provides a model-internal reliability signal that is typically absent from standard steering-vector pipelines and many test-time adaptation methods. This enables a conservative, label-free fallback: when reconstruction error is high, we revert to the zero-shot prediction, preventing negative steering under distribution shift. An important future direction is to adapt the SAE per instance at test time via lightweight self-reconstruction to reduce error, improve the salient features, and therefore steering effectiveness.
Finally, our upper-bound analyses indicate substantial headroom when task-relevant sparse features can be selected reliably. This highlights a central gap: the SAE objective surfaces features that are salient for reconstruction, but the features most useful for downstream classification can differ and are task-dependent. In our current setup, the same SAE features are used independent of the task, and we uniformly amplify the active features. Bridging reconstruction saliency and task saliency motivates future work on task-aligned sparse steering, i.e., selecting and amplifying the subset of SAE features that most directly support the target decision. Appendix K takes an initial step in this direction with Prototype-Alignment Sparse Steering, inspired by prototype theory (Rosch, 1973), suggesting that incorporating weak or indirect task structure during representation learning can better align sparse features with downstream needs.
6 Conclusion
We introduced Visual Sparse Steering (VS2), a lightweight, label-free domain adaptation method that steers vision foundation models using sparse features learned by top-$k$ Sparse Autoencoders, without updating model weights or requiring contrastive anchor examples. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves over zero-shot CLIP by 3.45%–4.12%, 0.93%–1.08%, and 1.50%–1.84%, respectively, while remaining forward-only at inference. To study the potential benefits of selective feature amplification, we presented VS2++ as a retrieval-based upper-bound analysis that uses retrieved neighbors to form positive/negative groups and construct a contrastive steering direction in SAE feature space. With oracle positives/negatives, VS2++ yields absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate. When the retrieved sets are noisy, performance can degrade, highlighting the need for more robust feature selection under weak supervision. We additionally observe that VS2 and VS2++ can disproportionately benefit visually or taxonomically proximate classes, improving per-class accuracy by up to 25% and 38%, respectively. Finally, VS2 provides a reconstruction-error-based reliability diagnostic that can disable steering and fall back to the baseline under a strict threshold, reducing the risk of harmful interventions at the cost of occasionally foregoing gains.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
Appendix A Decoding the Sparse Latent Space: Insights from SAEs
Our proposed methods achieve significant classification performance gains, primarily due to the contribution of Sparse Autoencoders (SAEs). We hypothesize that SAEs identify meaningful sparse features, which in turn guide the steering mechanisms. To validate this assumption, we conduct both quantitative and qualitative evaluations to assess the significance of the learned features.
A.1 Quantitative Evaluation of Feature Significance
To evaluate the role of sparse features, we manipulate the top-$k$ most active latent features extracted via Sparse Autoencoders. This experiment examines whether these features are critical for classification and how model predictions change under different modifications. We explore the following manipulation settings:
  1. 1.
    Zeroing-Out ($\gamma=0$): We set the top-$k$ most active features to zero before applying the steering vector. This removes their influence while preserving the remaining latent structure.
  2. 2.
    Negation ($\gamma=-1$): We invert the sign of the top-$k$ features before applying the steering vector, effectively pushing the representation in the opposite direction. This tests whether these dimensions encode class-discriminative information.
Table 7 reports the zero-shot classification accuracy of $\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ using ViT-B/32 after applying these transformations to the Sparse Steering vector intervention. We observe that negating or zeroing the top-$k$ most important features consistently degrades performance across all datasets. This result confirms that the features learned by the Sparse Autoencoders are essential for classification.
Table 7: Effect of manipulating top-$k$ sparse codes. Zero-shot accuracy (%) drops sharply when dominant sparse features are zeroed ($\gamma=0$) or negated ($\gamma=-1$), confirming their importance.
Modification CIFAR-100 Tiny-IN
$\text{CLIP}_{\text{ZS}}$ $61.07$ $56.64$
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ $62.69$ $39.49$
+ Zero-out ($\gamma=0$) $1.71$ $16.11$
+ Negate ($\gamma=-1$) $0.06$ $0.82$
A.2 Qualitative Evaluation of Features Significance
We qualitatively investigate the features learned in the sparse latent representations of the top-$k$ activations in the Sparse Autoencoder (SAE). Specifically, we assess the learned features by analyzing feature activations for each input and identifying the inputs that exhibit the highest activations for a given feature. Unlike mechanistic interpretability in the language domain, where an LLM can be used to assign semantic labels to a feature by summarizing its highly activated inputs, the vision domain lacks an equivalent automated labeling process.
To avoid reliance on human qualitative evaluation, we leverage annotated datasets where each image is associated with predefined attributes. For the qualitative evaluation of feature significance, we use the CUB dataset, which provides rich concept annotations for each image, enabling a structured assessment of the learned representations. Specifically, we investigate whether we can identify specific latent features with the highest concept coverage among their top-k most activated images. Concept coverage refers to how consistently a specific interpretable concept (e.g., an identifiable object category, attribute, or semantic idea) appears across a set of highly activated examples for a given SAE dimension. The intuition is that if a specific concept frequently emerges among the top-activating images for a particular feature, that feature is strongly associated with that concept.
For feature 511 as shown in Figure 3 (left), the activated images exhibit consistent semantic characteristics, including a gray upper part and a white underpart. Notably, this feature predominantly activates for images from similar but different classes, specifically different types of Gulls, such as Western, California, Herring, and Slaty-Backed Gulls. We observe that the top-activating images for these classes share all concepts in common which is not the case with dimension 3067. For feature 3067, as shown in Figure 3 (right) we observe that the top-activating images share common visual attributes, such as a white-colored throat and black eyes. Through human qualitative evaluation, we find that features in the sparse latent space capture meaningful visual concepts, grouping semantically similar images together, either from the same class or across different classes, as long as they share underlying conceptual similarities.
Figure 3: Concept coverage analysis of learned sparse latent features in the Sparse Autoencoder (SAE). Each subfigure illustrates the top-activating images for two different SAE dimensions, highlighting the consistency of shared visual concepts among the highest-activated examples. The analysis demonstrates how sparse features capture meaningful semantic attributes, grouping semantically similar images either within or across classes. Left: Top-activating images for feature 511. The images predominantly belong to different classes of Gulls (e.g., Western, California, Herring, and Slaty-Backed Gulls), yet they share consistent visual characteristics such as a gray upper part and a white underpart. This suggests that feature 511 captures a semantically meaningful concept spanning multiple related categories. Right: Top-activating images for feature 3067. The images share distinct visual attributes, including a white-colored throat and black eyes. However, unlike feature 511, these images belong to more diverse categories, indicating that this latent dimension captures a broader concept that generalizes across different classes.
Table 8: Top-10 most activated classes for SAE Feature 511 (left) and Feature 3067 (right). Feature 511 strongly aligns with gull-like classes, while Feature 3067 captures a cross-class head/throat attribute.
Class (511) Activation Class (3067) Activation 066.Western_Gull 96.7% (29/30) 082.Ringed_Kingfisher 60.0% (18/30) 060.Glaucous_winged_Gull 96.6% (28/29) 083.White_breasted_Kingfisher 40.0% (12/30) 061.Heermann_Gull 93.3% (28/30) 079.Belted_Kingfisher 30.0% (9/30) 059.California_Gull 90.0% (27/30) 041.Scissor_tailed_Flycatcher 23.3% (7/30) 062.Herring_Gull 83.3% (25/30) 080.Green_Kingfisher 16.7% (5/30) 063.Ivory_Gull 76.7% (23/30) 002.Laysan_Albatross 13.3% (4/30) 087.Mallard 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 147.Least_Tern 76.7% (23/30) 025.Pelagic_Cormorant 13.3% (4/30) 064.Ring_billed_Gull 70.0% (21/30) 044.Frigatebird 10.0% (3/30) 084.Red_legged_Kittiwake 73.9% (17/23) 049.Boat_tailed_Grackle 10.0% (3/30)
Class-Conditional Feature Activation.
To provide additional evidence that SAE dimensions capture discriminative visual concepts, we report the class-conditional activation frequencies of two example features: 511 and 3067. Specifically, Table 8 shows the percentage of samples in the top-10 activating classes where each feature is present.
What we observe is that feature 511 fires very often in gull-like classes. It also fires Larus-type seabirds, i.e., terns, kittiwakes. We observe that it also fires moderately on similar seabirds (Horned Puffin: 53.3%, White Pelican: 40%, Brown Pelican: 16.7%, Red-breasted Merganser: 10%). Additionally, we observe that Feature 511 activates near 0% on unrelated classes such as buntings, warblers, and hummingbirds. In contrast, Feature 3067 captures a cross-category visual attribute rather than a taxonomic grouping. Specifically, it consistently activates on species that share a distinctive head-and-throat pattern (e.g., white throat with darker eye region), even when those species are taxonomically unrelated. This indicates that the SAE possibly learns attribute-level discriminative concepts, complementing the class-specific features exemplified by Feature 511.
Discussion.
We note that our goal is not to claim human interpretability of SAE features. As shown in recent mechanistic interpretability work (Chanin et al., ), human-aligned features are often split across multiple SAE components (feature absorption), a known behavior in sparse dictionary learning. Nonetheless, our contribution lies in demonstrating that SAEs can go beyond interpretability, serving as effective mechanisms for performance improvement via sparse steering. While some features are interpretable, our main result is that sparse features enable controllable improvements in downstream classification tasks.
Appendix B Sensitivity to Sparse Amplification ($\gamma$) and Steering Magnitude ($\lambda$)
We analyze how the two key hyperparameters in VS2 (i) the sparse-feature amplification $\gamma$, and (ii) the steering vector scale $\lambda$ affect downstream accuracy. In the absence of contrastive supervision, these parameters govern how strongly we amplify sparse activations and how far the embedding is shifted in feature space. We sweep both values across a grid on three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet using ViT-B/32 backbone.
Figure 4 shows that all tested combinations of $(\lambda,\gamma)$ outperform the zero-shot baseline, though to varying degrees. Each dataset exhibits a diagonal band of near-optimal settings where $\lambda\cdot\gamma\in[2,3]$ tends to yield peak accuracy. For example, CIFAR-100 peaks at $\lambda^{*}=2.1$ and $\gamma^{*}=1.5$. Beyond this band, increasing $\lambda$ or $\gamma$ causes performance to degrade likely due to over-amplification of sparse features and/or embedding distortion. The consistent contour patterns across datasets suggest that VS2 is robust to moderate variation in its hyperparameters and that good settings generalize well across domains.
(a) CIFAR-100
(b) CUB-200
(c) Tiny-ImageNet
Figure 4: Sensitivity of VS2 to sparse amplification $\gamma$ and steering magnitude $\lambda$. All three datasets show a range of near-optimal combinations (warm colours), typically when $\lambda\cdot\gamma\in[2,3]$. Accuracy degrades if either parameter becomes too large.
Appendix C Training top-$k$ Sparse Autoencoders
We follow CLIP (Radford et al., 2021b) with a ViT-B/32 and ViT-B/16 vision backbones, intercepting the output of each encoder layer for the $CLS$ token. Specifically, we train top-$k$ SAEs on the $CLS$ embeddings for each chosen layer. We use $k=64$ and $k=256$ as the maximum active features within the latent space for ViT-B/32 and ViT-B/16 respectively, and we set a “dead feature” threshold of 100 i.e., any feature seldom activated is pruned. We also use an expansion factor of 4 relative to the input embedding dimension, resulting in 3,072 latent units. Training largely follows Gao et al. (2024a) and uses  EleutherAI (2024), with a linear learning-rate schedule and warmup from $5\times 10^{-4}$.
We monitor reconstruction quality with the fraction of variance unexplained (FVU; lower is better), defined as
$\text{FVU}\;=\;\frac{\lVert\mathbf{X}-\hat{\mathbf{X}}\rVert_{F}^{2}}{\lVert\mathbf{X}-\bar{\mathbf{X}}\rVert_{F}^{2}},$ (5)
where $\mathbf{X}\in\mathbb{R}^{B\times d}$ is a batch of CLS embeddings, $\hat{\mathbf{X}}$ denotes their SAE reconstructions, and $\bar{\mathbf{X}}$ is the batch mean (so the denominator equals the total variance). FVU is the complement of the coefficient of determination ($1-R^{2}$); an FVU of 0 indicates perfect reconstruction, while an FVU of $1$ corresponds to predicting only the mean. In Table 9, we report FVU results for ViT-B/32 with $k=64$ and ViT-B/16 with $k=256$ across all three datasets.
Table 9: Fraction of variance unexplained (FVU; lower is better). Each SAE has 4× expansion; sparsity $k=64$ for ViT-B/32 and $k=256$ for ViT-B/16.
Dataset ViT-B/32 (k=64) ViT-B/16 (k=256)
CIFAR-100 $0.2812$ $0.1166$
CUB-200 $0.2487$ $0.1653$
Tiny-ImageNet $0.5060$ $0.3018$
Appendix D Sensitivity in Retrieval-Augmented Generation (RAG)
When an external image corpus is available, VS2 can be extended using a Retrieval-Augmented Generation (RAG) pipeline. We use DINOv2 (Oquab et al., 2023) to retrieve top-$k$ images most similar to the input query and compute an enhanced embedding by combining the query with the retrieved set:
$\mathbf{E}=\alpha\mathbf{q}+(1-\alpha)\sum_{j=1}^{k}w_{j}\mathbf{r}_{j},$
where $\mathbf{q}$ is the query embedding, $\mathbf{r}_{j}$ the $j$-th retrieved embedding, and $w_{j}$ the normalized similarity weight:
$w_{j}=\frac{s_{j}}{\sum_{i=1}^{k}s_{i}},\quad\text{where }s_{j}=\text{sim}(\mathbf{q},\mathbf{r}_{j}).$
The parameter $\alpha\in[0,1]$ controls the trade-off between using the original query and the retrieved set. We sweep over values of $\alpha$ and $k$ to assess their impact on zero-shot classification performance. Figure 5 shows results for CIFAR-100, and Tiny-ImageNet. CIFAR-100 and Tiny-ImageNet display a similar trend: larger $\alpha$ (i.e., more reliance on the query) typically degrades performance. On Tiny-ImageNet, setting $\alpha=0$, completely ignoring the input query and relying purely on retrieved features, yields the best result. Across all datasets, large $k$ eventually hurts performance, confirming that RAG benefits from focused rather than broad context. The trade-off parameter $\alpha$ and the retrieval depth $k$ play dataset-dependent roles in RAG-enhanced pipelines. For fine-grained domains, fewer, high-confidence neighbors and a low $\alpha$ might work best; for noisy domains, more reliance on the retrieval images might be more beneficial.
(a) CIFAR-100
(b) Tiny-ImageNet
Figure 5: RAG sensitivity to $\alpha$ and top-$k$. Accuracy varies with the weight $\alpha$ on the original query and the number $k$ of retrieved images. Larger $k$ often introduces noise; smaller $\alpha$ performs better on cluttered datasets.
Appendix E Steering Across Layers of the CLS Token
Our default method reconstructs and applies steering to the CLS token embedding at the final layer of the vision encoder. To study depth effects, we evaluate steering the CLS token across multiple final layers of the transformer. We use ViT-B/16 on CIFAR-100 with expansion factor 4 and 128 sparse activations. Table 10 reports top-1 accuracy when steering the last 1, 2, or 3 layers. Steering only the final layer yields the best performance. Adding earlier layers causes a progressive drop in accuracy, falling below the zero-shot baseline when steering the last three layers.
Table 10: Effect of steering depth on CLS. Top-1 accuracy (%) on CIFAR-100 (ViT-B/16) as the number of steered final layers increases.
Steered Layers Accuracy (%)
12 68.08
11 + 12 65.72
10 + 11 + 12 59.36
Applying steering at multiple layers likely introduces compounding perturbations that propagate forward, making later representations harder to align with the classifier. CLS steering is most effective at the final layer. Future work could explore coordinated multi-layer steering to avoid error accumulation.
Appendix F Ablation Study of Expansion factor and top-$k$
In Table 11, we present the downstream task accuracy of CLIP ViT-B/32 using various values of expansion factor and $k$. Across a 4 $\times$ range in width and an 8 $\times$ range in sparsity, top-1 accuracy fluctuates by less than one percentage point evidence that VS2’s performance is largely insensitive to the precise SAE capacity–sparsity trade-off. Additionally, in Table 18, we present the average cosine similarities of different steering vectors coming from various configurations of SAEs in terms of expansion factor and top-$k$.
Table 11: VS2 accuracy as a function of SAE width (expansion factor) and sparsity (top-$k$). Numbers are top-1 / top-5 (%). The best result is boldfaced; every other configuration is within of the optimum, highlighting the method’s robustness to architectural choices.
SAE configuration Top-1 $\uparrow$ Top-5 $\uparrow$
4 $\times$, kk=128 64.61 87.95
8 $\times$, kk=128 64.56 87.76
4 $\times$, kk=64 64.54 87.79
16 $\times$, kk=64 64.54 87.78
8 $\times$, kk=64 64.52 87.96
10 $\times$, kk=128 64.43 87.81
16 $\times$, kk=512 64.42 87.71
8 $\times$, kk=512 64.40 87.91
4 $\times$, kk=256 64.34 87.62
8 $\times$, kk=256 64.29 87.80
16 $\times$, kk=256 64.28 87.75
4 $\times$, kk=512 64.12 87.87
16 $\times$, kk=128 64.10 87.79
Appendix G Visual Sparse Steering Pseudocode
For reproducibility purposes, in Algorithm 1, we provide the pseudocode for the baselines and VS2 used in the analysis of 4.1.
Algorithm 1 SAE_Steering – SAE-based CLS token modification
1:function SAE_Steering($\mathbf{h},\mathrm{SAE},k,\gamma,\lambda_{\Delta z},\text{mode}$) $\triangleright$ $\mathbf{h}$: CLS token; SAE: sparse autoencoder; $k$: sparsity level
2:  $\mathbf{a},\mathbf{idx}\leftarrow\mathrm{SAE.select\_topk}(\mathrm{SAE.pre\_acts}(\mathbf{h}),k)$
3:  $\mathbf{z}_{\text{base}}\leftarrow\mathrm{SAE.decode}(\mathbf{a},\mathbf{idx})$
4:  $\mathbf{z}_{\text{boost}}\leftarrow\mathrm{SAE.decode}(\gamma\cdot\mathbf{a},\mathbf{idx})$
5:  if $\text{mode}=\texttt{"reconstruction"}$ then
6:   return $\mathbf{z}_{\text{base}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}}$
7:  else if $\text{mode}=\texttt{"amplified"}$ then
8:   return $\mathbf{z}_{\text{boost}}$ $\triangleright$ Variant: $\text{CLIP}_{\text{REC}}^{\text{F}+\gamma}$
9:  else if $\text{mode}=\texttt{"steering"}$ then
10:   return $\mathbf{h}+\lambda_{\Delta z}\cdot(\mathbf{z}_{\text{boost}}-\mathbf{z}_{\text{base}})$ $\triangleright$ VS2
11:  end if
12:end function
Appendix H Comparison with other Baseline Methods
We compare VS2 against SpLiCE (Bhalla et al., 2024), using the official implementation provided by the authors. For all three datasets, we report performance using SpLiCE with an external vocabulary of 10,000 LAION-based concepts and an $\ell_{1}$ regularization weight of 0.25, following their best reported configuration. Despite relying on no external vocabulary, VS2 consistently outperforms SpLiCE across all benchmarks highlighting the strength of sparse concept steering even in the absence of external lexical resources.
Table 12: Zero-shot top-1 accuracy (%) on CIFAR-100.
Method ViT-B/32 ViT-B/16
$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 )
$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 )
$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 )
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 )
SpLiCE (Bhalla et al., 2024) 55.57 ( –5.50 ) 58.29 ( –5.67 )
VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 )
Appendix I Effect of Top-$N$ Retrieved Neighbors
To assess the influence of retrieval size on performance, we conduct an ablation over the number of retrieved neighbors $N$ used for latent aggregation. We use a fixed ViT-B/16 (Patch-16) backbone and vary $N\in\{10,25,50,100\}$ across three datasets: CIFAR-100, CUB-200, and Tiny-ImageNet. As shown in Figure 6, both top-1 and top-5 classification accuracy steadily increase with $N$ on CIFAR-100 and Tiny-ImageNet, with diminishing returns beyond $N=50$. In contrast, performance on CUB-200 remains largely flat or slightly degrades, suggesting that retrieving more neighbors in fine-grained datasets can introduce noise due to overly similar but semantically irrelevant examples. These results indicate that the utility of neighbor-based aggregation is dataset-dependent: general object recognition tasks may benefit from larger $N$, while fine-grained classification may require more careful control of retrieval scope.
Figure 6: Top-1 and Top-5 accuracy as a function of number of retrieved neighbors NN, using ViT-B/16 (Patch-16) across datasets. Larger NN improves general classification but may degrade performance in fine-grained settings.
Appendix J Per-class Performance Analysis on CIFAR-100
In Table 17, we present results for the classes most affected by steering CLIP ViT-B/32 with VS2 and VS2++. We highlight both positive and negative shifts in accuracy to identify the most impacted categories. Unlike the average performance reported in the main experiments, this analysis shows that the top-5 class-level gains are generally larger in magnitude than the corresponding losses. Although the accuracy drops for misclassified categories are smaller in magnitude, they are non-negligible. As future work, it would be valuable to explore when steering should be applied or withheld to further maximize overall performance gains.
Appendix K Prototype-Aware Sparse Steering Vectors
Our core methods, VS2 and its retrieval-augmented variant VS2++, enhance zero-shot classification by steering CLIP embeddings along directions identified by a top-$k$ Sparse Autoencoder (SAE). These directions correspond to latent features that, ideally, align with class-discriminative concepts. Steering in these directions upweights what the model has learned to be important during reconstruction. This raises a central hypothesis: the reconstruction task itself is to some extent sufficient to uncover features that are also relevant for classification. In other words, there is a meaningful overlap between features that are important for reconstructing the CLS token and those that are predictive for the downstream task. In this section, drawing inspiration from prototype theory (Rosch, 1973), we investigate whether incorporating prototype information during SAE training can better align the features important for reconstruction with those that are critical for downstream classification.
The limited improvements observed on fine-grained datasets like CUB-200 and Tiny-ImageNet suggest that the challenge lies not just in identifying sparse features, but in uncovering the correct ones. This shifts the central question from “What are the most important sparse features to select?” to a deeper inquiry: “Can the reconstruction objective alone reliably capture features that are most useful for classification and if not, how can task-relevant information be effectively incorporated?”
Table 13: Benchmarking Visual Sparse Steering: Zero-Shot Accuracy (%) with and Without External Data on CIFAR-100, CUB-200, and Tiny Imagenet using ViT-B/32 and ViT-B/16.
CIFAR-100 CUB-200 Tiny-IN
Method ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16 ViT-B/32 ViT-B/16
Zero-shot (no retrieval)
$\text{CLIP}_{\text{ZS}}$ 61.07 ( 0 ) 63.96 ( 0 ) 51.76 ( 0 ) 55.06 ( 0 ) 56.64 ( 0 ) 61.08 ( 0 )
$\text{SAE}_{\text{REC}}^{\text{A}}$ 58.01 ( –3.06 ) 64.05 ( +0.09 ) 47.45 ( -4.31 ) 51.81 ( -3.25 ) 30.56 ( –26.08 ) 52.96 ( -8.12 )
$\text{SAE}_{\text{REC}}^{\text{F}}$ 58.22 ( –2.85 ) 63.42 ( –0.54 ) 48.08 ( -3.68 ) 51.43 ( -3.63 ) 36.33 ( –20.31 ) 54.84 ( -6.24 )
$\text{SAE}_{\text{REC}}^{\text{F}+\gamma}$ 62.69 ( +1.62 ) 66.81 ( +2.85 ) 49.41 ( -2.35 ) 53.28 ( -1.78 ) 39.49 ( –17.15 ) 58.81 ( -2.27 )
VS2 (ours) 64.52 ( +3.45 ) 68.08 ( +4.12 ) 52.69 ( +0.93 ) 56.14 ( +1.08 ) 58.14 ( +1.50 ) 62.92 ( +1.84 )
VS2 + PASS (ours) 70.64 ( +9.57 ) 68.23 ( +4.27 ) 52.97 ( +1.21 ) 56.63 ( +1.57 ) 57.94 ( +1.30 ) 62.98 ( +1.90 )
Oracle steering with known prototypes.
Following prototype theory, we collect for every class the ten images that ViT-B/32 CLIP classifies with the highest confidence; these serve as oracle prototypes. Given the true class label $y$, we build a prototype steering vector by averaging their latent sparse features. VS2 using oracle prototypes can lift CLIP to 97.5% on CIFAR-100, 91.04% on CUB-200, and 90.1% on Tiny-ImageNet. This confirms the existence of discriminative sparse directions. In Appendix L we further examine the discriminative ability (measured by orthogonality) of these steering vectors. Generally, these prototype vectors have a low cosine similarity, yet a non-negligible tail reveals strongly overlapping directions between visually or semantically close categories.
Prototype-aligned SAE (PASS).
Above, we constructed steering vectors using the pretrained SAE features of oracle prototypes. Now, we consider whether prototypes can be used to learn new, more informative SAE features. We assume during SAE training that we have access to class labels. Then, to the SAE loss we add a regularization term which encourages SAE features to be close to their class mean. That is, for a training sample $i$ with latent sparse feature  $\mathbf{z}_{i}$ and class mean $\bar{\mathbf{z}}_{\text{class}(i)}$ we minimize
$\mathcal{L}=\mathcal{L}_{\text{recon}}+w_{\text{aux}}\;\bigl\lVert\mathbf{z}_{i}-\bar{\mathbf{z}}_{\text{class}(i)}\bigr\rVert_{2}^{2},$ (6)
where $w_{\text{aux}}$ controls the strength of the prototype-alignment term relative to the reconstruction loss and is set to 0.8. We refer to the resulting steering method as PASS (Prototype-Aligned Sparse Steering). Although PASS uses class labels during SAE training, it remains fully test-time unsupervised. Empirically, in Table 13, we observe that PASS outperforms VS2 across all datasets, with particularly substantial gains on CIFAR-100. However, this improvement comes at the cost of requiring labels for each training sample during SAE training. Gains are modest on CUB-200 and Tiny-ImageNet, and we thus hypothesize that classes which share many features require richer or multi-prototype guidance which is an intriguing avenue for future work.
Table 14: Trade-off between reconstruction fidelity and prototype alignment. Increasing $w_{\text{aux}}$ improves classification accuracy but degrades FVU reconstruction.
$w_{\text{aux}}$ FVU $\downarrow$ Accuracy (%) $\uparrow$
0.1 0.3437 68.80
0.5 0.4887 70.40
1.0 0.5393 70.61
2.0 0.5702 70.76
Reconstruction vs. Alignment Trade-off
Sparse Autoencoders (SAEs) trained for reconstruction can also be optimized to align their latent features with class-level prototypes. However, this introduces a trade-off between two competing objectives: fidelity of reconstruction and discriminative alignment.
To investigate this trade-off, we introduce a weighting coefficient $w_{\text{aux}}$ that controls the strength of prototype alignment relative to the reconstruction loss. As $w_{\text{aux}}$ increases, alignment is encouraged more strongly. Table 14 reports the resulting changes in reconstruction loss (measured by FVU) and top-1 classification accuracy on CIFAR-100 using ViT-B/16, with 128 sparse latents and expansion factor 4.
We observe that increasing $w_{\text{aux}}$ consistently improves classification performance, up to a point, even though it introduces more reconstruction error. This aligns with our hypothesis: while exact input reconstruction encourages general feature coverage, alignment with class prototypes promotes discriminative feature extraction. These results highlight the flexibility of SAE-based steering to balance interpretability and performance depending on downstream objectives.
Appendix L Are exemplar-derived directions really distinct?
A desirable property of class–specific steering vectors is orthogonality: pushing an embedding toward class A should not simultaneously raise its score for class B. Using the oracle prototypes as described in Appendix K, we compute for every CIFAR-100 class a prototype steering vector and measure the pair-wise cosine similarity at layer 11 of the ViT-B/32 encoder. Most pairs have low similarity (mean=0.23), yet a non-negligible tail reveals strongly overlapping directions. Table 15 lists the ten highest-overlap pairs.
Table 15: Top-10 most overlapping prototype steering directions on CIFAR-100. High cosine similarity indicates that the two classes share visual attributes that the SAE encodes along nearly the same latent axis.
Rank Class 1 Class 2 Cosine $\uparrow$
1 beetle cockroach $0.91$
2 mouse shrew $0.89$
3 dolphin shark $0.84$
4 otter seal $0.84$
5 dolphin whale $0.84$
6 possum raccoon $0.84$
7 snake worm $0.83$
8 oak tree willow tree $0.83$
9 ray shark $0.81$
10 bowl cup $0.80$
Table 16: Training-time compute and parameter cost per sample for various adaptation strategies on the CLIP ViT-B/32 vision encoder.
Method Train FLOPs / sample (fwd+bwd) GFLOPs Trainable Params
SAE (cached activations) 14,164,992 0.014 4,722,432
SAE (non-cached; incl. extraction) 8,743,645,440 8.744 4,722,432
LoRA (rank = 16) 26,365,388,544 26.365 589,824
Full fine-tuning (vision encoder) 26,188,441,344 26.188 87,456,000
These high-overlap pairs are semantically plausible confusions (e.g. beetle vs. cockroach or dolphin vs. whale), confirming that exemplar steering directions tend to align for visually or taxonomically proximate classes. In downstream applications, a simple orthogonalization step may help reduce feature overlap between sparse directions. Investigating principled ways to encourage orthogonality during SAE training is a promising direction for future work.
Table 17: Top-5 class gains and losses on CIFAR-100. Green = absolute gain; Red = absolute loss relative to ZS baseline.
Class ZS VS2 $\uparrow$($\Delta$) tractor $0.55$ 0.80 $\uparrow$(+0.25) forest $0.35$ 0.58 $\uparrow$(+0.23) man $0.53$ 0.74 $\uparrow$(+0.21) bus $0.51$ 0.68 $\uparrow$(+0.17) snake $0.61$ 0.76 $\uparrow$(+0.15)
(a) VS2 – Top-5 Gains
Class ZS VS2 ($\Delta$) aquarium_fish $0.82$ 0.61 (-0.21) beetle $0.64$ 0.49 (-0.15) sweet_pepper $0.70$ 0.58 (-0.12) tulip $0.78$ 0.71 (-0.07) maple_tree $0.57$ 0.50 (-0.07)
(b) VS2 – Top-5 Losses
Class ZS VS2++ $\uparrow$($\Delta$) spider $0.48$ 0.91 $\uparrow$(+0.43) caterpillar $0.27$ 0.68 $\uparrow$(+0.41) possum $0.24$ 0.65 $\uparrow$(+0.41) tractor $0.55$ 0.96 $\uparrow$(+0.41) tiger $0.45$ 0.84 $\uparrow$(+0.39)
(c) VS2++ – Top-5 Gains
Class ZS VS2++ ($\Delta$) girl $0.72$ 0.65 (-0.07) maple_tree $0.57$ 0.51 (-0.06) porcupine $0.18$ 0.13 (-0.05) ray $0.06$ 0.02 (-0.04) mouse $0.18$ 0.15 (-0.03)
(d) VS2++ – Top-5 Losses
Appendix M Computational Overhead: Additional Details
Training-time FLOPs.
To quantify the cost of training the SAE versus alternative adaptation methods, Table 16 reports per-sample training FLOPs (forward + backward) and the number of trainable parameters for the CLIP ViT-B/32 vision encoder. Even when activation extraction is included, SAE training remains more than $3\times$ cheaper than LoRA or full fine-tuning. With cached activations, as in our VS2 setup, the cost drops to just 0.014 GFLOPs per sample.
TPT FLOP accounting.
For completeness, we outline how the $16.6\times$ test-time overhead factor for TPT (Shu et al., 2022) is obtained. A single CLIP ViT-B/32 forward pass on a $224\times 224$ image with 100 text prompts costs $8.7$ GFLOPs for the vision encoder and $581.8$ GFLOPs for the text encoder, for a total of approximately $0.59$ TFLOPs. In our implementation, TPT performs $63$ augmented views per image and runs $4$ optimization steps; each step requires a full CLIP forward pass over all views and a backward pass through the text encoder. This yields about $9.217$ TFLOPs of adaptation compute per test image, followed by one final forward-only evaluation of $0.591$ TFLOPs, for a total of roughly $9.807$ TFLOPs. Dividing by the $0.59$ TFLOPs of plain CLIP gives the reported $\approx 16.6\times$ test-time compute overhead.
Table 18: Cosine similarity ($\uparrow$) between steering vectors learned by different SAE capacities. Rows/columns are ordered by expansion factor and sparsity (e × expansion, $k$ active).
SAE cfg. 16×/64 16×/512 16×/128 16×/256 8×/128 8×/512 8×/256 4×/256 4×/512 4×/128 10×/128 4×/64 8×/64 16×/64 1.000 0.497 0.577 0.705 0.726 0.635 0.680 0.640 0.591 0.681 0.536 0.701 0.716 16×/512 0.497 1.000 0.161 0.399 0.358 0.817 0.634 0.618 0.611 0.584 0.136 0.316 0.282 16×/128 0.577 0.161 1.000 0.900 0.865 0.273 0.298 0.277 0.265 0.303 0.952 0.887 0.678 16×/256 0.705 0.399 0.900 1.000 0.956 0.527 0.572 0.495 0.410 0.542 0.920 0.936 0.744 8×/128 0.726 0.358 0.865 0.956 1.000 0.543 0.650 0.575 0.483 0.634 0.896 0.959 0.829 8×/512 0.635 0.817 0.273 0.527 0.543 1.000 0.861 0.811 0.771 0.796 0.243 0.483 0.506 8×/256 0.680 0.634 0.298 0.572 0.650 0.861 1.000 0.889 0.771 0.912 0.289 0.557 0.624 4×/256 0.640 0.618 0.277 0.495 0.575 0.811 0.889 1.000 0.873 0.941 0.240 0.532 0.600 4×/512 0.591 0.611 0.265 0.410 0.483 0.771 0.771 0.873 1.000 0.821 0.191 0.461 0.549 4×/128 0.681 0.584 0.303 0.542 0.634 0.796 0.912 0.941 0.821 1.000 0.284 0.594 0.655 10×/128 0.536 0.136 0.952 0.920 0.896 0.243 0.289 0.240 0.191 0.284 1.000 0.910 0.678 4×/64 0.701 0.316 0.887 0.936 0.959 0.483 0.557 0.532 0.461 0.594 0.910 1.000 0.814 8×/64 0.716 0.282 0.678 0.744 0.829 0.506 0.624 0.600 0.549 0.655 0.678 0.814 1.000